### v2

# The Challenge of Using LLMs to Simulate Human Behavior: A Causal Inference Perspective

Abstract
Large Language Models (LLMs) have shown impressive potential to simulate human behavior. We identify a fundamental challenge in using them to simulate experiments: when LLM-simulated subjects are blind to the experimental design (as is standard practice with human subjects), variations in treatment systematically affect unspecified variables that should remain constant, violating the unconfoundedness assumption. Using demand estimation as a context and an actual experiment as a benchmark, we show this can lead to implausible results. While confounding may in principle be addressed by controlling for covariates, this can compromise ecological validity in the context of LLM simulations: controlled covariates become artificially salient in the simulated decision process, which introduces focalism. This trade-off between unconfoundedness and ecological validity is usually absent in traditional experimental design and represents a unique challenge in LLM simulations. We formalize this challenge theoretically, showing it stems from ambiguous prompting strategies, and hence cannot be fully addressed by improving training data or by fine-tuning. Alternative approaches that unblind the experimental design to the LLM show promise. Our findings suggest that effectively leveraging LLMs for experimental simulations requires fundamentally rethinking established experimental design practices rather than simply adapting protocols developed for human subjects.
1 Introduction
The rise of large language models (LLMs) like GPT has sparked interest across disciplines (including marketing, computer science, economics, psychology and political science) in leveraging these tools to simulate how humans would respond to questions or stimuli in different contexts (Aher et al.,, 2023; Argyle et al.,, 2023; Arora et al.,, 2024; Brand et al.,, 2023; Dillion et al.,, 2023; Goli and Singh,, 2024; Horton,, 2023; Park et al.,, 2023). If these LLM-simulated experiments accurately approximate real human behavior, the implications for both academics and practitioners are substantial. Academics could use LLMs for pilot experiments to pinpoint stimuli with significant impact, thus improving the efficiency of theory development and experimental design. In fact, several review articles have already been written about this recent yet fast-growing literature (Bail,, 2024; Hutson and Mastin,, 2023; Sarstedt et al.,, 2024; Simmons and Hare,, 2023). Firms could leverage these realistic simulations to explore different ideas and strategies, thereby improving customer insight and product development and optimizing marketing mix variables. Accordingly, in the recent past we have witnessed a large influx of firms offering services leveraging LLMs for customer insights (e.g., Synthetic Users, Outset AI, Nexxt, Voxpopme, Evidenza, Expected Parrot, Meaningful, xPolls, Ipsos, and CivicSync).
Many practical applications of LLM simulations involve exploring counterfactual scenarios where one factor is altered while others remain constant, hence requiring causal estimates of treatment effects. For example, what would happen to sales if prices were increased, how would consumers respond if a message were framed differently, or how would market dynamics shift with the introduction of a new product? While the success of using LLMs to simulate counterfactual depends on the size and quality of their training or validation data that can correct for potential bias (Ludwig et al.,, 2024; Wang et al.,, 2024), the success also critically depends on the design of the simulations.
Although extensive literature on experimental design with human subjects has established best practices and universal standards—such as random assignment and blinding—it remains unclear what constitutes best practices in designing LLM simulations and how these might differ from traditional experimental design principles. Establishing LLM experimental design principles is crucial for enhancing the reliability and validity of LLM simulation results, thereby making them more valuable for both researchers and practitioners. Without clear guidelines on which experimental designs to adopt or avoid, the vast flexibility and sensitivity in LLM simulations can lead to widely varying outcomes (Carlson and Burbano,, 2024; Gao et al.,, 2024; Ludwig et al.,, 2024; Mohammadi,, 2024). When different simulation approaches yield conflicting results, researchers lack a basis for determining how much weight to place on each finding. By developing standardized design principles, researchers can assign greater weight to well-designed simulations, enhancing the reliability and usefulness of the results. Developing such design principle is essential for advancing the field and ensuring that LLM simulations effectively contribute to our understanding of human behavior.
Many LLM simulations to date have relied on blinded experimental designs, where simulated participants are not aware of the experimental conditions or hypotheses being tested (Argyle et al.,, 2023; Brand et al.,, 2023; Horton,, 2023). Blinding [^1] Also sometimes referred to as “masking” (Association et al.,, 2019). has many benefits in human subject research: it reduces demand effects where participants modify their behavior based on what they believe researchers want to see and often helps maintain ecological validity by presenting scenarios that more closely mirror real-world decision contexts. Following this convention, researchers typically construct LLM prompts that present a single scenario without revealing its experimental nature, similar to how human participants would experience the treatment condition in a traditional between-subjects design.
Although blinding is considered a best practice in traditional experiments, we show that it can have unintended consequences in LLM simulations which may no longer justify its benefits. Unlike human subjects and contexts that exist independently of the experiment, LLM-simulated subjects and contexts are ”created on the fly” based on the prompt. When researchers make LLM-simulated “subjects” blind to experimental conditions, variations in the treatment can systematically affect unspecified variables that should remain constant, violating the unconfoundedness assumption typically required for valid causal inference.
We use demand estimation as an example to empirically illustrate this challenge. We conduct experiments simulating consumer demand at different price points. We find that altering the product’s price not only affects the purchase outcomes but also systematically influences other variables that should not be influenced by the treatment—such as pre-treatment variables (e.g., demographic characteristics and customer purchase history) and contextual factors (e.g., price of competing products). This unintended variation in unspecified covariates violates the unconfoundedness assumptions, and we show that it also generates implausible demand curves when compared against an actual experiment with a representative sample of human participants.
A tempting solution is specifying more variables in the prompts to address this confounding issue. However, we find this approach may sacrifice ecological validity because including more covariates can make the scenarios presented to the LLM less reflective of real-world decision-making contexts. For instance, when competing product information is explicitly mentioned in prompts, these variables become artificially salient in the LLM’s decision process, unlike typical consumers who may not always consider such information during routine purchases. This focalism issue creates a fundamental tension between controlling for confounding factors and maintaining ecological validity - a tradeoff that typically does not exist in traditional experiments where covariates can be measured and controlled for during data analysis without making them artificially salient to participants during data generation.
We present a theoretical framework illustrating that confoundedness in blind LLM simulations stems from prompts being ambiguous. To overcome this fundamental limitation and improve the simulation results, researchers should consider relaxing the blinding requirement and informing the LLM of the experimental design. This approach is not subject to the same confounding issues and holds promise for better generalizable performance as LLMs continue to advance. To complement this approach, an opportunity for future research is to develop LLMs that can better interpret and understand experimental designs as well as natural language descriptions of counterfactual scenarios.
More broadly, we aim to contribute to the growing literature on LLM simulations by providing a starting point for establishing best practices in their design. By highlighting the trade-offs between blinding, unconfoundedness, and ecological validity, we offer insights that can help researchers and practitioners design more reliable and valid LLM-based experiments. Our work underscores the importance of adjusting traditional experimental design principles to the unique characteristics of LLMs, ensuring that these powerful tools can be effectively harnessed to simulate and understand human behavior.
The rest of the paper is organized as follows. Section 2 demonstrates the unintended consequences of blinding empirically using demand estimation as an example. Section 3 explores the standard solution of controlling for covariates and highlights its unique challenges in the context of LLM simulations. Section 4 proves an impossibility theorem demonstrating this fundamental issue will persist for future models as long as ambiguous prompting strategies are used, and examines its implications for techniques like fine-tuning and RAG. Section 5 discusses promising directions for making prompts unambiguous through unblinding, and Section 6 concludes.
2 Unintended Confounding in Blind LLM Simulations
The principle of blinding is a cornerstone of experimental design that consists in withholding information from the participants, such as hiding the design of the experiment or the hypotheses being tested. It is widely used in between-subject experiments with human participants, and is also sometimes implemented in within-subject experiments. [^2] For example, in a blind taste test comparing Coca-Cola and Pepsi, participants evaluate both beverages (within-subject design) but remain blind to which drink is which, the hypothesis about brand preferences, the tasting order, and the total number of samples they’ll evaluate. Blinding helps reduce demand effects and maintains ecological validity by presenting scenarios that mirror real-world decision contexts. This section uses GPT4, demand estimation and one common LLM simulation design to illustrate how blinding creates unique challenges in LLM simulations that are absent in human experiments, and Section 4 generalizes the findings to other versions of LLMs, other applications, and other simulation designs.
One common way to implement blinding in LLM simulations is to conduct between-prompt experiments, where each prompt represents a single experimental condition without revealing the broader experimental context (Argyle et al.,, 2023; Brand et al.,, 2023; Horton,, 2023). Blinded between-prompt LLM experiments have been used to mimic both between-subject and within-subject experiments with human participants. For example, to mimic a between-subject experiment that studies price sensitivity, researchers might present different prices to the LLM in separate prompts, with each prompt describing only the current price scenario without mentioning the experimental nature of the price variation. To mimic a within-subject conjoint analysis study, researchers might present different product profiles or choice sets to the LLM in separate prompts, with each prompt describing only the current product profile or choice set.
While random assignment in traditional human experiments ensures pre-treatment variables are balanced across conditions and contextual variables remain constant (even when unobserved), LLM-simulated participants and environments are generated dynamically. When researchers leave details unspecified in blind LLM prompts, the model must make (implicit or explicit) assumptions about these unspecified variables, which can systematically vary with the treatment and introduce confounding factors.
We illustrate these issues using demand estimation, a classical problem in economics and marketing, where researchers are often interested in understanding how changes in a product’s price affect demand while holding everything else constant. We select the top 40 Consumer Packaged Goods (CPG) categories according to DellaVigna and Gentzkow, (2019), and for each category we pick one of the top selling products according to Walmart.com. We record the regular price of each product on Walmart.com around April, 2024. Table A.1 in the Appendix documents this set of products. For each product, we vary the price from 0 to 200% of the regular price, in 20% increments. This gives us 11 price points to test for each product ($0\%\}$ , 20 % , ..100 % , ..200 % } of the regular price). Throughout the paper, for each prompt template and for each {category,product} pair described in Table A.1, we vary the price systematically across these 11 price points. We refer to “relative price” as the price relative to the regular price ($\{-1,-0.8,...,0,...,0.8,1\}$ where 00 is the regular price). We use the Open AI API and we “pin” the model to gpt-4o-mini-2024-07-18 (the latest available at the time of running the experiment) to make sure that the exact same model is used for all our simulations. We adjust the “System” part of the prompt to increase the probability of receiving a usable answer from the model (the “System” portion of the prompt, which can be manipulated on the API and which is set to “You are a helpful assistant” on ChatGPT, provides meta-instructions to GPT). Unless noted otherwise, for each prompt that corresponds to a {category, product, price} combination, we set the temperature to 1 and generate 50 draws to calculate the average response.
First, we elicit unspecified variables from the LLM. For each {category, product, price} combination, we submit a prompt that asks GPT to fill in the blank contextual or pre-treatment variables given the randomized price (Kurita et al.,, 2019; Nozza et al.,, 2022; Zhang et al.,, 2020). Prompt 2 gives an example of such template that asks the past price given the current randomized price.
Prompt 1: Unintended correlation between price and past price System: You, AI, are a customer. Your task is to fill in the blanks __. Return the completed information in comma-separated values, without any extra text. User: Please consider the following product category: {category}. Suppose you are in a grocery store, and you see the following product in that category: {product}. The last time you purchased this product, it was priced at \$          [a number with up to 2 decimal points]. The product is currently priced at: {price}. Return example 1: XX.XX
Figure 1 demonstrates positive correlations between the price of the focal product and several variables, including the price that the customer paid last time, the price of the competing product, and the expiration days of the product. The x-axis is price relative to the baseline price, allowing us to aggregate across products.333 Prompts B and B in Appendix B document the prompts used to elicit the correlations between randomized price and competing price/expiration date. Appendix C presents additional analysis demonstrating the broader scope of these unintended correlations across multiple variables. While these correlations are plausible in observational data, they are undesirable in an experiment where researchers are interested in understanding what happens when estimating the treatment effect of changing the price of the focal product while holding everything else constant.
Figure 1: Unintended correlation between price and past price, competing price, and expiration days
Next, we explore whether these unintended correlations are associated with confounding and biased estimates of relevant quantities when the goal is estimating causal treatment effects. To that end, we run an actual survey on human subjects, use a similar prompt in an LLM simulation, and compare the results. The survey with human participants (pre-registered, see https://aspredicted.org/pqys-st6w.pdf) was developed using Qualtrics. Each respondent answered one purchase intention question per product, with prices randomly drawn for each product and the order of products randomized across respondent. [^4] We acknowledge that one difference between the study run with humans vs. GPT is that each human answered 40 questions, while on GPT a different independent call to the API was made for each question (following standard practice). Following our pre-registration, we collected data from 1,000 respondents, and excluded 9 respondents according to pre-registered exclusion criteria (based on attention filters at the beginning of the survey). The respondents were recruited from Prolific, where we paid a premium to obtain a representative sample of the US population (based on census data). We then use Prompt 2 to simulate potential purchase decisions at different price points (where again {category,product} and {price} were varied according to our experimental design, mirroring the experiment with humans), and for each corresponding prompt, we used gpt-4o-mini-2024-07-18, with temperature set to 1, to generate 50 draws to calculate the average purchase probability at each {category,product,price} combination. Figure 2 shows the average demand curves obtained from GPT and from human participants. While the curves elicited from human participants follow a classic downward-sloping demand curve, the GPT-simulated curves are different from the curves elicited from human participants and follow an inverted-U shape. Although it is understandable that it is challenging for GPT to predict demand when price equals 0 as such scenario is probably not well represented in its training data, the curve remains inverted-U shape even if we remove the first point corresponding to a price of zero.
Prompt 2: Ask Purchase System: You, AI, are a customer. Your task is to fill in the blanks  . Return the completed information in comma-separated values, without any extra text. User: Please consider the following product category: {category}. Suppose you are in a grocery store, and you see the following product in that category: {product}. The product is currently priced at \${Price}. Would you or would you not purchase the product?   [“purchase” or “not purchase”] Return example: purchase
Figure 2: Demand curve elicited from humans vs. LLM
These findings are not surprising in hindsight: in real observational data, the price of the focal product should be informative of the price of the competing products, especially if we do not know anything about the context, the type of store, and how the price for a product is set. An LLM trained on such data would also naturally make such association. This association can be realistic and useful if researchers intend to use LLMs to simulate competing prices based on the focal product price. However, this same association is problematic when researchers intend to estimate the causal impact of changing the focal price, while holding everything else constant, because it leads to endogeneity. One common approach to addressing such endogeneity is to control for additional covariates. The next section explores the benefits and novel challenges associated with this approach.
3 Potential Solution: Controlling for Covariates
When working with real-world data, a common approach for causal identification is to control for covariates. If researchers have controlled for all confounding covariates that affect both the treatment and outcome, the conditional independence assumption, or ignorability, is satisfied. If sufficient variation remains after controlling for covariates, then researchers can identify the causal effect of interest. In LLM-simulated experiments, a natural way to control for covariates is adding detailed information about individuals and their environment in the prompt. These covariates can range from demographic information, location, preferences, budget, etc. We discuss other approaches including RAG and fine-tuning in the Section 4.3.
Argyle et al., (2023) introduced an approach for controlling for such covariates. They advocate for running studies on “silicon samples” of digital personas, which in their case were described by demographic variables and political beliefs that were drawn to align with nationally representative samples. This approach has been picked up by the industry and commercialized by several firms (e.g., Synthetic Users, Evidenza). Figure 3 shows that on average, controlling for covariates, such as demographic variables does indeed improve the results in our case compared to not specifying any detail. However, it does not fully resolve the issue. Appendix D describes the detailed prompts and covariates used, which involve the creating of 500 personas for each {product,category} pair.
Figure 3: Demand curves elicited by LLM after controlling for demographic variables
An important consideration is the stage at which covariates are controlled for. When working with human data, empirical researchers often control for product characteristics and rich user history to address endogeneity during data analysis, after the data is already collected. Hence analyzing data does not change the original data generating process that already took place. In contrast, adding covariates in the prompt in LLM simulation occurs at the data generation stage. Consequently, the additional covariates can change the factors considered by a simulated individual when making decisions.
For example, in the context of demand estimation, one can include contextual information in the prompt such as the price of competing products in the category, which can be used as a proxy for the type of stores that the customer is visiting and the competitive environment that the focal product is facing. However, while adding contextual variables such as competing prices in the LLM simulation may prevent the LLM from using the price treatment to infer these contextual variables, adding them may also lead the LLM-simulated customer to put a heavier emphasis on these factors when making decisions. Figure 4 shows that the demand curve (elicited by simulating the choices of 500 personas) is no longer smooth after fixing such contextual variables, [^5] We fix the competing price to be the same as the regular price of the focal product. and the simulated customers’ purchase decisions follow a step function: customers will purchase if and only if the price is lower than or equal to the competing price. Appendix E describes the detailed prompts. [^6] One possible way to fix the non-smoothness of the demand curve is to simulate a distribution of contextual variables, each with difference past and competing prices. However, this is no longer useful for a particular context, e.g., a particular store, because the simulated context deviates from the targeted context. In a real-world scenario, we should expect that even in the absence of heterogeneity in context (e.g., in a given store in a given day), there should still exists heterogeneity in consumers, such that the demand curve would be smoother.
Figure 4: Purchase probability of Coca-Cola when controlling for demographic variables and competing price
Such phenomenon in which decision makers place more emphasis on certain aspects of a choice set or scenario than they would in real-world decisions is known in the consumer behavior literature as focalism (Schkade and Kahneman,, 1998; Wilson et al.,, 2000), and constitutes a threat to ecological validity. For example, in the context of conjoint analysis, Bedi and Reibstein, (2021) argue that when minor attributes are presented together with major ones, they tend to receive undue attention from respondents (due to focalism), leading to upward biased estimates of the importance of these minor features. While it has been documented that LLM is prone to focalism (Cheng et al.,, 2023), it is important to emphasize that focalism is not an issue that is unique to LLM experiments.
While traditional experiments with humans also run the risk of failing to simulate realistic scenarios, the nature of this risk differs. When running experiments with humans, using a blind design with realistic scenarios typically ensures ecological validity without introducing concerns of confoundedness. On the other hand, with LLM simulations, a blind design raises confoundedness concerns. Adressing confoundedness by adding covariates requires adding covariates in the data generation stage of the study, which raises yet another concern, focalism, which in turn reduces ecological validity. In other words, LLM simulations pose a unique dilemma to researchers. Improving ecological validity may require leaving unspecified details in the prompt, which runs the risk of introducing confoundedness. On the other hand, including extensive detail may reduce confounding, but at the expense of ecological validity due to focalism.
4 Theoretical Framework
So far we have demonstrated one empirical example with one specific version of an LLM. A natural question is whether the issue we have uncovered applies generally to all LLMs. This section presents a theoretical framework demonstrating that the challenges identified in this paper stems from ambiguous prompts rather than fundamental limitations of the models themselves. Therefore, the solution to this specific challenge lies not only in developing more advanced LLMs, but also in developing clearer prompting strategies.
4.1 Blinding Generates Ambiguous Prompts
To demonstrate how ambiguity arises in LLM prompts, we begin by examining a simple prompt $p$ under a blinded simulation design that illustrates the fundamental problem:
Prompt 3: A simple ambiguous prompt Please consider the following product category: {category}. Suppose you are in a grocery store, and you see the following product in that category: {product}. The product is priced at: {price}. Would you or would you not purchase the product?
The design being blind, incomplete information is provided about the distribution from which the various variables in the prompt are drawn. As a result, this prompt is ambiguous. It allows for at least two interpretations:
  • $q_{1}$ (interventional): what is the customer’s purchase decision when the price is set to be $\{price\}$?
  • $q_{2}$ (observational): what is the customer’s purchase decision when the price happens to be $\{price\}$?
The answers to these two questions are likely to be different. Using the do-operator (Pearl,, 2009) to distinguish mere observations and active intervention, and using $Y$ to represent $Purchase$ and $D$ to represent Price, the answer to $q_{1}$ can be represented as $a_{1}:P(Y|do(D))$ while the answer to $q_{2}$ can be represented as $a_{2}:P(Y|D)$.
Consider a true data-generating process (DGP) described by a linear structural equation:
$Y=\beta_{M}D+U$
where $\beta_{M}$ represents the causal effect of $D$ on $Y$ through a set of realistic mechanism $M$, and $U$ represents unobserved factors. In a well-designed experiment, we have $Cov(U,D)=0$, making it possible to identify $\beta_{M}$ and estimate $P(Y|do(D))$ from data on $(Y,D)$ alone. But under a blinded design with an ambiguous prompt, the LLM does not know it should mimic an experimental environment. That is, the LLM does not have enough information to draw $U$ from the appropriate distribution (literally or figuratively - we do not attempt to model how the LLM makes inference but rather present a conceptual model). If the model is mostly trained on observational data, it is mostly likely going to mimic observational data patterns, where $Cov(U,D)\neq 0$. As a result, the LLM’s responses will reflect confounded relationships rather than the intended causal effects. Figure 5 summarizes two possible DGPs mimicked by the LLM under this simple ambiguous prompt.
$D$$M$$Y$$U$(a) DGP for Experimental $D$$M$$Y$$U$(b) DGP for Observational Data
Figure 5: Possible DGPs mimicked by LLM simulations under ambiguous simple prompts
As discussed in Section 3, one might attempt to address this issue by adding more variables $X$ to the prompt. In principle, if $Cov(U,D|X)=0$, then controlling for $X$ can help restore identification. However, simply specifying covariates in the LLM prompt does more than condition on them statistically. Including $X$ may cause the simulation to alter the underlying decision-making mechanisms $M$ and shift toward a new set of mechanisms $M^{\prime}$ and produce an environment that differs substantially from the original scenario – the estimated parameter $\beta_{M^{\prime}}$ may have limited ecological validity, since it no longer reflects the realistic mechanisms $M$ that were originally of interest (as we illustrated in Figure 4). Figure 6 summarizes two possible DGPs mimicked by the LLM under this more complicated ambiguous prompt.
[…]
Figure 6: Potential DGPs mimicked by LLM simulations when controlling for additional variables
These graphs reveal a deeper challenge: adding variables to the prompt can actually increase rather than decrease ambiguity from a causal perspective. To illustrate this multiplication of possible interpretations in practice, consider the following detailed example:
Prompt 4: More Detailed Scenario with Ambiguous Treatment User: You are a customer who has recently read an article about how sugar consumption affects health. You are now in a store and you see Regular Coke (39g sugar per 12 oz, 12 pack cans) priced at \$4.99, with a soda tax of \$0.12/oz. Would you or would you not purchase the product?
This prompt, while detailed, is even more ambiguous. It permits more valid causal interpretations:
  1. 1.
[…]
    “What happens when we experimentally set price to \$4.99, holding fixed that customers are health-aware?”
  2. 2.
[…]
  3. 3.
[…]
4.2 Impossibility Theorem for Ambiguous Prompting Strategies
Having demonstrated how ambiguity can arise in practice, we now formalize these insights into a theoretical framework that explains why such ambiguity poses fundamental challenges for LLM evaluation and development.
Definition 4.1 (Functional Representation of a Large Language Model).
A large language model may be represented by:
$f:\mathcal{P}\rightarrow\Delta(\mathcal{Y}),$
where $\mathcal{P}$ is the set of all possible prompts, $\mathcal{Y}$ is the set of all possible outputs or responses, and $\Delta(\mathcal{Y})$ represents the set of all probability distributions over $\mathcal{Y}$. For any given prompt $p\in\mathcal{P}$, the model $f$ outputs a probability distribution $f(p)$ over potential responses $y\in\mathcal{Y}$.
[…]
Definition 4.2 (Prompting Strategy).
A prompting strategy is a function:
$s:\mathcal{Q}\rightarrow\mathcal{P},$
that maps each question $q\in\mathcal{Q}$ to a specific prompt $p\in\mathcal{P}$.
While prompting strategies can be refined empirically using validation datasets, developing a theoretical framework is essential for understanding when and why these strategies will generalize beyond the validation set. When ground truth data is available, researchers can experiment with different prompts and even fine-tune LLMs to improve performance on specific questions. However, without theoretical guidance, such empirical refinements may not yield strategies that reliably generalize to novel scenarios. This raises a fundamental question: what characteristics make a prompting strategy robust and reliable for out-of-distribution inference? We identify a particular class of prompting strategies to avoid:
Definition 4.3 (Ambiguous Prompting Strategy).
An ambiguous prompting strategy occurs when there exists distinct questions $q_{1}$ and $q_{2}$ that have different correct answers, but are mapped to the same prompt $p$: $s(q_{1})=s(q_{2})=p$ for some prompt $p\in\mathcal{P}$.
Such ambiguous prompting strategies should be avoided because no LLM models, no matter how advanced they are, can answer all questions correctly given such prompting strategy:
Theorem 4.1.
[Impossibility Theorem for Ambiguous Prompting Strategies] For a given ambiguous prompting strategy, there exists no LLM $f$ that can correctly answer all questions correctly.
Proof.
Assume that $s(q_{1})=s(q_{2})=p$ for two distinct questions $q_{1}$ and $q_{2}$ that have different correct answers $a_{1}$ and $a_{2}$, with $a_{1}\neq a_{2}$. If the model can correctly answer $q_{1}$, such that $f(s(q_{1}))=f(p)=a_{1}$, then the model cannot answer $q_{2}$ because $f(s(q_{2}))=f(p)=a_{1}\neq a_{2}$. ∎
4.3 Implications for model development using fine-tuning or RAG
Theorem 4.1 reveals how ambiguous prompting strategies create limitations when developing and deploying LLMs, even with techniques like fine-tuning or retrieval-augmented generation (RAG). Suppose developers leverage additional training data $(q_{1},a_{1})$ with a prompting strategy where $p=s(q_{1})$. If this prompt is ambiguous such that $p=s(q_{2})$ for some other question $q_{2}$ with correct answer $a_{2}\neq a_{1}$, then any fine-tuning that optimizes the model’s accuracy on $q_{1}$ necessarily sacrifices its accuracy on $q_{2}$. This trade-off is especially problematic given uncertainty about future applications - users may interpret the same prompts in different ways which developers cannot anticipate during model development.
Consider a concrete example where developers have access to an experimental dataset that randomizes Coca-Cola’s price while keeping Pepsi’s price at business-as-usual levels. If developers fine-tune an LLM using a blinded design that simply describes purchase decisions at different Coca-Cola prices without revealing their experimental nature, they create an ambiguous prompt that permits both experimental and observational interpretations. While the fine-tuned model may excel at predicting purchases in scenarios matching the training data ($q_{1}$), this capability is less valuable since the experimental data already directly answers such questions. The more important question is whether the model can learn the causal relationship between price and purchases, and can generalize to answer novel questions ($q_{2}$) - like what happens when Pepsi’s price changes while Coca-Cola’s price remains at business-as-usual levels, or when both prices increase simultaneously. Such generalization proves challenging theoretically.
First, under ambiguous prompts, fine-tuning based on experimental data provides no guarantee that the model will learn causal relationships. Suppose that the data generating process for the experimental dataset follows $Y=\beta_{M}D+U$, where $Y$ represents the purchase decision, $D$ represents the price of Coca-Cola, and $U$ represents everything else. The fine-tuned model may implicitly learn a regression coefficient $\widehat{b}=\frac{Cov(Y,D)}{Var(D)}=\beta_{M}+\frac{Cov(U,D)}{Var(D)}$. However, since the fine-tuned model builds on the base model, it can maintain a prior that $\widehat{\frac{Cov(U,D)}{Var(D)}}=\rho\neq 0$ and hence will not interpret the regression coefficient $\widehat{b}$ as causal. After accounting for this prior, the model may even incorrectly assume that the causal effect is $\widehat{\beta}_{M}=\widehat{b}-\rho$.
Furthermore, even if the model somehow updates its assumptions correctly to handle one specific type of question ($q_{1}$), it may fail to generalize to other closely related but distinct questions ($q_{2}$) where different assumptions are needed. For example, suppose the model learns from experimental data where Coca-Cola’s price was randomized while Pepsi’s price remained business-as-usual. Through fine-tuning, the model might correctly learn to treat Coca-Cola’s price as independent of other unobserved market factors when analyzing that specific scenario ($q_{1}$). However, when asked to predict what happens when Pepsi’s price is set to another level while Coca-Cola’s price remain business-as-usual ($q_{2}$), the model might incorrectly maintain this assumption of independence, when in fact Coca-Cola’s business-as-usual price would likely be correlated with other market factors in this new scenario. This limitation is avoidable if the experimental nature of the data is explicitly communicated to the model during both training and evaluation through less ambiguous prompts.
While the above example focuses on a case with one experimental dataset, the challenge persists even when using diverse experimental datasets for fine-tuning or retrieval-augmented generation (RAG). With RAG, this challenge manifests at both the development and deployment stages. During model development, the reference datasets should be annotated with metadata specifying which variables were randomized treatments versus correlated covariates. During retrieval, prompts should unambiguously specify the causal question to ensure the model draws from experiments with matching causal structures rather than just similar surface features. Without such explicit specification of the causal framework, neither fine-tuning nor RAG can reliably learn to distinguish between different types of variation in the data.
To summarize, our theoretical framework proves that no general-purpose model can accurately simulate all counterfactual scenarios under an ambiguous prompting strategy. This statement should not be interpreted as suggesting that fine-tuning or RAG are not valuable tools - these techniques can substantially enhance model performance by incorporating high-quality training data and reducing various forms of bias. Rather, it demonstrates that while better training data and advanced techniques are important for improving simulation accuracy, they are complementary to, rather than substitutes for, clear and unambiguous prompt design. While obtaining better data for validation, fine-tuning, or bias-correction can be resource-intensive, implementing unambiguous simulation designs may be achieved at no additional cost. The optimal approach combines both strategies: leveraging advanced training techniques while ensuring an unambiguous prompting strategy.
5 Unblinding as a Potential Solution
Our theoretical framework points toward a promising direction: designing unambiguous prompts by explicitly communicating the experimental design in the prompt, which by definition entails unblinding. Here we discuss several possible simulation designs that leverage unblinding.
One approach maintains the between-prompt structure where each prompt simulates one experimental condition, but explicitly reveals the full experimental design in each prompt. This allows generating multiple responses to form a distribution for each condition while maintaining the independence of observations across prompts. Another possibility is to use a within-prompt design where multiple experimental conditions are presented together in a single prompt with full information on the experimental design. This approach is more efficient computationally since it obtains responses for multiple conditions in one call to the model. However, showing multiple conditions together may induce artificial consistency in the responses. [^8] It is worth noting that within-prompt designs are not necessarily unblinded - a prompt could present multiple conditions (e.g., different prices) without explaining how those conditions were selected or distributed. Therefore, merely switching to within-prompt experiments is not sufficient to resolve prompt ambiguity. An even simpler approach is an unblind within-prompt aggregate design that directly asks the LLM to predict the aggregate results of the full experiment. This represents the greatest departure from traditional experimental protocols with human subjects.
Comparing the relative performance of these different approaches is beyond the scope of this paper. Unlike our analysis of blinding - where we demonstrated fundamental limitations that will persist across model generations - the effectiveness of different unblinded designs may vary substantially across models and contexts. The relative performance one may find now may no longer hold in the future, especially if future models are trained or fine-tuned based on these prompts. Due to this reason, this section only focuses on demonstrating the feasibility of the unblinded design through the simplest approach of directly eliciting aggregate predictions of the results of the experiment.
For this demonstration, we directly ask the LLM to predict aggregate experimental results across randomly assigned prices. Figure 7 shows that while this strategy’s predictions still deviate from actual consumer behavior, it does outperform the other potential solutions tested in this paper. Appendix F documents the detailed prompts.
Figure 7: Unblinded design asking directly for aggregate predictions
This approach is by no means perfect. In addition to the fact that the result still deviates from the ground truth, Figure F.1 shows that the results are also sensitive to the experimental range provided in the prompt: when the demand curve is built by combining two experiments that vary prices from [0%, 100%] and from [100%, 200%] of the regular price respectively, rather than a single experiment with prices varying from [0%, 200%], the estimates change. There are several possible reasons for this. One is that this LLM is unable to correctly process experimental information. Another possibility is that the existing prompt still has some remaining ambiguity. For example, the LLM may make assumptions about the setting, such as the type of stores and the customers, based on the set of prices included in the experiment. Appendix F conducts additional analysis which suggests that this sensitivity to the experimental range is partly driven by the current LLMs not being able to correctly process experimental design information, and shows that newer LLMs demonstrate improved (though still imperfect) consistency across different experimental ranges. Therefore, unlike the fundamental limitations of blinding that future LLMs may not easily address, sensitivity to experimental design information could potentially be resolved through model improvements, making this a promising area for future research.
Figure 8: Unblinded design asking directly for aggregate predictions: price range comparison
6 Conclusion
This paper identifies a fundamental challenge in using LLMs to simulate experiments with human participants. Unlike human participants who exist independently of experiments, LLM-simulated individuals and environments are generated based on prompts. While LLMs’ ability to capture associations from training data is generally beneficial, it creates complications in experimental settings where changing one variable can unintentionally affect other unspecified variables meant to remain constant. This confounding makes it difficult to accurately simulate counterfactual human behavior. More generally and fundamentally, making the LLM blind to the experimental design leads to ambiguous prompts. Given LLMs are trained primarily on observational data, ambiguous prompts tend to favor correlational rather than causal observations, making variables unspecified in the prompt susceptible to the intended intervention. Because these challenges stem from ambiguous prompting strategies rather than model limitations, our theoretical framework suggests that it is not possible to fully resolve it through improved training data.
Our findings have important practical implications for different stakeholders in the LLM ecosystem. For researchers evaluating LLM performance, our results caution against using blinded designs when assessing simulation capabilities. Poor performance in such settings may reflect prompt ambiguity rather than model limitations, making it impossible to properly diagnose model capabilities. For academic researchers studying LLMs, our work highlights the need to establish consensus on unambiguous prompting strategies. Without such consensus, different research teams may map distinct questions to identical prompts, leading to conflicting interpretations that no model can simultaneously satisfy. Perhaps most importantly, our findings suggest avoiding ambiguous prompts during the development and fine-tuning of LLMs. When prompts permit multiple interpretations, improving performance on one interpretation may degrade performance on others - an unnecessary trade-off that could be avoided through clearer prompting strategies, especially given uncertainty about future applications.
We explored several potential solutions to this challenge. Adding detailed context to control for confounders showed promise but tends to introduce focalism which could invalidate results. A more promising approach consists in unblinded experimental designs that explicitly communicate randomization. This approach showed potential for reducing ambiguity, though its effectiveness may vary across models and contexts. Therefore, developing LLMs that can better follow the instruction to simulate counterfactual human behavior given unambiguous prompts can be a crucial direction for future research.
Our paper highlights a broader insight of using LLMs: rather than simply applying human subject protocols to LLMs, researchers need to develop new experimental practices tailored to LLMs’ unique properties. The field of LLM simulations is still emerging, and best practices remain to be established. Our work provides a theoretical foundation for understanding one of the key challenges and evaluating potential solutions. We hope this framework will guide future research in developing more reliable methods for leveraging LLMs in experimental research, while remaining mindful of their fundamental differences from human subjects.
Appendix A Category, Product and Price Information
Table A.1: Categories, products and regular prices
Category Product Price (\$)
Fruit Juice Capri Sun Variety Pack with Fruit Punch, Strawberry Kiwi & Pacific Cooler Juice Box Pouches, 30 ct Box, 6 fl oz Pouches 9.43
Fruit Drinks Kool Aid Jammers Variety Pack with Tropical Punch, Grape & Cherry Kids Drink 0% Juice Box Pouches, 30 Ct Box, 6 fl oz Pouches 7.27
Baby Milk and Milk Flavoring Horizon Organic Shelf-Stable Whole Milk Boxes, 8 oz., 12 Pack 13.98
Soup Maruchan Ramen Noodle Chicken Flavor Soup, 3 Oz, 12 Count Shelf Stable Package 9.97
Cat Food - Wet Type Purina Fancy Feast Chicken Feast Classic Grain Free Wet Cat Food Pate - 3 oz. Can 0.88
Pet Supplies - Dog Food Purina Dog Chow Complete, Dry Dog Food for Adult Dogs High Protein, Real Chicken, 44 lb Bag 29.17
Snacks - Potato Chips Lay’s Classic Potato Snack Chips, Party Size, 13 oz Bag 5.44
Snacks - Tortilla Chips Doritos Nacho Cheese Tortilla Snack Chips, Party Size, 14.5 oz Bag 5.94
Cereal - Ready to Eat Cinnamon Toast Crunch Breakfast Cereal, Crispy Cinnamon Cereal, Family Size, 18.8 oz 4.93
Cookies Little Debbie Oatmeal Creme Pies, 12 ct, 16.2 oz 2.68
Ground and Whole Bean Coffee Folgers Classic Roast Ground Coffee, Medium Roast, 40.3-Ounce Canister 13.24
Soft Drinks - Carbonated Coca-Cola Soda Pop, 12 fl oz, 12 Pack Cans 8.26
Bottled Water OZARKA Brand 100% Natural Spring Water, 16.9-ounce plastic bottles (Pack of 35) 19.96
Candy - Chocolate Hershey’s Milk Chocolate Candy, Bars 1.55 oz, 6 Count 6.48
Candy - Non-Chocolate HARIBO Goldbears Original Gummy Bears, 28.8oz Stand Up Bag 6.48
Soft Drinks - Low Calorie Coca-Cola Zero Sugar Soda Pop, 16.9 fl oz, 6 Pack Cans 5.18
Frozen Italian Entrees Smart Ones Three Cheese Ziti Marinara Frozen Meal, 9 Oz Box 2.26
Frozen Foods Great Value All Natural Chicken Wing Sections, 4 lb (Frozen) 12.98
Ice Cream Haagen Dazs Coffee Ice Cream, Gluten Free, Kosher, 14.0 oz 4.18
Frozen Novelties Pop-Ice Assorted Fruit Freezer Ice Pops, Gluten-Free Snack, 1.5 oz, 80 Count Fruit Pops 6.17
Lunchmeat - Sliced - Refrigerated Oscar Mayer Chopped Ham & Water product Deli Lunch Meat, 16 Oz Package 4.33
Frankfurters - Refrigerated Oscar Mayer Classic Uncured Beef Franks Hot Dogs, 10 ct Pack 3.94
Refrigerated Bacon Oscar Mayer Fully Cooked Original Bacon, 2.52 oz Box 4.27
Refrigerated Entrees John Soules Foods Chicken Breast Fajita Strips, Refrigerated, 16oz, 18g Protein per 3oz Serving Size 5.98
Dairy Products Land O Lakes Salted Stick Butter, 16 oz, 4 Sticks 5.28
Yogurt - Refrigerated Chobani Non-Fat Greek Yogurt, Vanilla Blended 32 oz, Plastic 5.58
Refrigerated Deli Meats Goya Cooked Ham 16 oz 29.99
Dairy - Milk - Refrigerated Great Value Milk Whole Vitamin D Gallon 3.92
Bakery - Fresh Cakes Little Debbie Zebra Cakes, 13 oz 2.68
Fresh Eggs Eggland’s Best Classic Extra Large White Eggs, 12 count 3.18
Fresh Fruit Fresh Raspberries, 12 oz Container 4.74
Beer Stella Artois Lager, 12 Pack, 11.2 fl oz Glass Bottles, 5% ABV, Domestic Beer 15.73
Light Beer (Low Calorie/Alcohol) Bud Light Beer, 24 Pack, 12 fl oz Aluminum Cans, 4.2% ABV, Domestic Lager 20.98
Detergents - Heavy Duty - Liquid Purex Liquid Laundry Detergent Plus OXI, Stain Defense Technology, 128 Fluid Ounces, 85 Wash Loads 9.97
Cleaning Supplies ARM & HAMMER Pure Baking Soda, For Baking, Cleaning & Deodorizing, 1 lb Box 1.54
Toilet Tissue Angel Soft Toilet Paper, 9 Mega Rolls, Soft and Strong Toilet Tissue 6.68
Paper Towels Bounty Select-a-Size Paper Towels, 12 Double Rolls, White 22.18
Batteries Duracell Coppertop AA Battery, Long Lasting Double A Batteries, 16 Pack 15.97
Pain Remedies - Headache Tylenol Extra Strength Caplets with 500 mg Acetaminophen, 100 Ct 10.97
Cold Remedies -Adult Equate Value Size Honey Lemon Cough Drops with Menthol, 160 Count 4.68
Note: The 40 categories were taken from DellaVigna and Gentzkow, (2019) and the top products and their regular prices from Walmart.com as of April, 2024. Due to the dynamic nature of retail pricing, actual prices may vary.
Appendix B Simple prompts for eliciting correlation
Prompts B and B document the prompts used to elicit the correlations between randomized price and competing price/expiration days shown in Figure 1.
Prompt 5: Correlation between price and competing price System: You, AI, are a customer. Your task is to fill in the blanks __. Return the completed information in comma-separated values, without any extra text. User: Please consider the following product category: {category}. Suppose you are in a grocery store, and you see the following product in that category: {product}. The product is currently priced at: {price}. The price of a similar competing product from a different brand is           [a number with up to 2 decimal points]. Return example 1: XX.XX
Prompt 6: Correlation between price and expiration days System: You, AI, are a customer. Your task is to fill in the blanks __. Return the completed information in comma-separated values, without any extra text. User: Please consider the following product category: {category}. Suppose you are in a grocery store, and you see the following product in that category: {product}. The product is currently priced at: {price}. The expiration date of the product is           [a whole number] days from now. Return example 1: 10
Appendix C Covariance of unspecified variables
This section demonstrates how prevalent variables covary with each other. We consider a more detailed prompt C.1 that elicits demographic, contextual, and pre-treatment variables.
Prompt 7: Correlation between multiple variables System: You, AI, are a customer. Your task is to fill in the blanks  . Return the completed information in comma-separated values, without any extra text. User: You are a consumer with the following characteristics: Age:   [a whole number] Gender:   Education level:   Household income:   [a whole number] Occupation:   Ethnicity:   Marital status:   Household size:   [a whole number] Number of children:   [a whole number] State of residence:   [state] Home ownership:   [e.g., “own,” “rent”] Please consider the following product category: {category}. Suppose you are in a grocery store, and you see the following product in that category: {product}. You have purchased this product   [e.g., “frequently,” “occasionally,” “rarely”] in the past. The last time you saw this product, it was priced at \$  [a number with up to 2 decimal points]. You have   [e.g., “a lot of,” “some,” “limited”] storage space at home. Your monthly grocery budget is   [a whole number]. The expiration date of the product is   [a whole number] days from now. The product is currently priced at \${Price}. A similar competing product from a different brand is priced at \$  [a number with up to 2 decimal points]. Would you or would you not purchase the product?   [“purchase” or “not purchase”] Return example: 35, female, bachelor’s degree, 50000, software engineer, Asian, married, 3, 1, CA, own, occasionally, 2.99, some, 800, 60, 3.49, purchase
Figure C.1: Eliciting correlation between relevant variables
Figure C.2 summarizes the covariance matrix between these variables. On one hand, this covariance matrix demonstrates that the LLM is capable of simulating many realistic correlations between covariates. For example, household income is positively correlated with budget. However, several covariates are correlated with the price change variable. Specifically, the focal price is correlated with product characteristics such as expiration days of the product, contextual variables such as competing prices, and pre-treatment variables such as purchase history. In an experiment with humans, these variables would either be held constant or be randomly distributed across treatment conditions in a way that is orthogonal to the treatment.
Figure C.2: Covariance matrix of demographic, contextual, and pre-treatment variables
Appendix D Simulation that controls for demographic variables
We use the following prompt template to first generate the distribution of demographic variables and then control for demographic variables. First, for each {category,product}, we leave the demographic variables as blank, and use the pinned version of gpt-4o-mini-2024-07-18 to generate 500 replications (personas) with a temperature of 1, giving us the joint distribution of demographic variables for that product. See Prompt D. To avoid imposing any prior assumption on Price, we leave {Price} as ”   [a number with up to 2 decimal points]” in this step. Then in the second step, we use the same prompt template but with the demographic variables filled in, i.e., all variables except the purchase decision filled in and the price being randomized. For each {category,product} and for each persona for that product, we vary the price, and elicit the purchase decision, while setting the temperature to 0, and then aggregate across the 500 personas to calculate the purchase probability. Then we aggregate across all products at the same relative price to construct the aggregate demand curve.
Prompt 8: Control for demographic variables System: You, AI, are a customer. Your task is to fill in the blanks  . Return the completed information in comma-separated values, without any extra text. User: You are a consumer with the following characteristics: Age:   [a whole number] Gender:   Education level:   Household income:   [a whole number] Occupation:   Ethnicity:   Marital status:   Household size:   [a whole number] Number of children:   [a whole number] State of residence:   [state] Home ownership:   [e.g., “own,” “rent”] Please consider the following product category: {category}. Suppose you are in a grocery store, and you see the following product in that category: {product}. The product is currently priced at   [a number with up to 2 decimal points]. Would you or would you not purchase {product}?   [“purchase” or “not purchase”] Return example: 35, female, bachelor’s degree, 50000, software engineer, Asian,married, 3, 1, CA, own, 3.99, purchase
Appendix E Simulation that controls for demographic variables + competing price
To provide a concrete example, we consider Coca-cola and asks what happens to the purchase probability when the price of a competing product is set at a certain level. We follow the exact same two-step procedure as Appendix D, except that we have fixed the price of a competing product to be the same as the regular price of the focal product when eliciting purchase decisions from the 500 personas.
Prompt 9: Control for demographic variables+ competing price System: You, AI, are a customer. Your task is to fill in the blanks  . Return the completed information in comma-separated values, without any extra text. User: You are a consumer with the following characteristics: Age: {age} Gender: {gender} Education level: {education} Household income: {income} Occupation: {occupation} Ethnicity: {ethnicity} Marital status: {marital status} Household size: {household size} Number of children: {number of children} State of residence: {state} Home ownership: {home ownership} Please consider the following product category: Soft Drinks - Carbonated. Suppose you are in a grocery store, and you see the following product in that category: Coca-Cola Soda Pop, 12 fl oz, 12 Pack Cans. The product is currently priced at \${Price}. A similar competing product from a different brand is priced at \$8.26. Would you or would you not purchase Coca-Cola Soda Pop, 12 fl oz, 12 Pack Cans?   [“purchase” or “not purchase”] Return example: purchase
This mimics the scenario of an actual store, where the price of a competing product is set uniformly at a certain level. Different customers, each with their own characteristics, will come to the store and make their purchase decision. While it is plausible that some customers may notice the price of the competing product and use it as a reference point to make purchase decisions, not all customers are attentive and may be aware of the price of all other competing products. Therefore, it is implausible to assume that the purchase will be $100\%$ ($0\%$) when the price of the focal product is lower than or equal to (greater than) the price of competing products.
Appendix F Direct simulation of the experimental results using unblinded design
Prompt F is a prompt template we use to predict the experimental results using an unblinded design, where information about the experiment is directly provided. For each {category,product}, we replace {RegularPrice} with the corresponding price listed in Table A.1.
For Figure 7, we replace the {Prices} with the set of prices that we experiment with, from 0% to 200% of regular price. For example, when the regular price is \$4.68. {Prices} are ”\$0.00, \$0.94, \$1.87, \$2.81, \$3.74, \$4.68, \$5.62, \$6.55, \$7.49, \$8.42, \$9.36”. In this basic version, we leave Additional Instruction 1 and 2 as empty strings “”. We set the temperature to 0 and use ’gpt-4o-mini-2024-07-18’.
For Figure 7, in addition to considering the {Prices} to be from 0% to 200%, we also try a version from 0% to 100%, labeled as ”Lower_Half_Price_Range” and a version from 100% to 200%, labeled as ”Upper_Half_Price_Range”. It shows that the simulation result is sensitive to the range of prices provided in the experiment.
To assess whether this sensitivity is driven by another form of prompt ambiguity, where LLMs may make unintended assumptions on the context given the set of experimental prices, in Figure F.1 we provide more explicit instructions by replacing {Additional Instruction 1} with ”Note that some elements of the experiments are chosen arbitrarily by the researchers and you should not rely on them to make your prediction.” and replacing {Additional Instruction 2} with ”Note that the Set of Prices to be experimented is chosen arbitrarily and you should not make your prediction based on this Set of Prices.” However, we find that the results are still sensitive, suggesting that the issue are driven by the current generation of LLMs that we are testing being unable to closely follow such instruction, rather than prompt ambiguity. To assess the potential for future models to make consistent prediction, we further test ’o1-mini-2024-09-12’ and ’o1-preview-2024-09-12’. Because they do not allow for ’System’ and only accept the ’User’ part of the prompt, we concatenate the ’System’ and ’User’ together as our input prompt. Figure F.1 shows that when using ’gpt-4o-mini-2024-07-18’, the inconsistency still remains substantial despite explicit instruction not to rely too much on the experimental range to make the prediction. However, ’o1-mini-2024-09-12’ and ’o1-preview-2024-09-12’ demonstrates increasingly decreased inconsistencies.
[…]
Figure F.1: Unblinded design with explicit instruction to disregard experimental range: Multiple consecutive model versions.

### v3

@@
 We identify a fundamental challenge in using them to simulate experiments: when LLM-simulated subjects are blind to the experimental design (as is standard practice with human subjects), variations in treatment systematically affect unspecified variables that should remain constant, violating the unconfoundedness assumption.
-Using demand estimation as a context and an actual experiment as a benchmark, we show this can lead to implausible results.
-While confounding may in principle be addressed by controlling for covariates, this can compromise ecological validity in the context of LLM simulations: controlled covariates become artificially salient in the simulated decision process, which introduces focalism.
-This trade-off between unconfoundedness and ecological validity is usually absent in traditional experimental design and represents a unique challenge in LLM simulations.
-We formalize this challenge theoretically, showing it stems from ambiguous prompting strategies, and hence cannot be fully addressed by improving training data or by fine-tuning.
-Alternative approaches that unblind the experimental design to the LLM show promise.
-Our findings suggest that effectively leveraging LLMs for experimental simulations requires fundamentally rethinking established experimental design practices rather than simply adapting protocols developed for human subjects.
+Using demand estimation as a context and an actual experiment with 40 different products as a benchmark, we show this can lead to implausible results.
+While confounding may in principle be addressed by controlling for covariates, this can compromise ecological validity in the context of LLM simulations: controlled covariates become artificially salient in the simulated decision process.
+We show formally that confoundness stems from ambiguous prompting strategies.
+Therefore, it can be addressed by developing unambiguous prompting strategies through unblinding, i.e., revealing the experiment design in LLM simulations.
+Our empirical results show that this strategy consistently enhances model performance across all tested models, including both out-of-box reasoning and non-reasoning models.
+We also show that it is a technique that complements fine-tuning: while fine-tuning can improve simulation performance, an unambiguous prompting strategy makes the predictions robust to the inclusion of irrelevant data in the fine-tuning process.
 1 Introduction
-The rise of large language models (LLMs) like GPT has sparked interest across disciplines (including marketing, computer science, economics, psychology and political science) in leveraging these tools to simulate how humans would respond to questions or stimuli in different contexts (Aher et al.,, 2023; Argyle et al.,, 2023; Arora et al.,, 2024; Brand et al.,, 2023; Dillion et al.,, 2023; Goli and Singh,, 2024; Horton,, 2023; Park et al.,, 2023).
+The rise of large language models (LLMs) like GPT has sparked interest across disciplines (including marketing, computer science, economics, psychology and political science) in leveraging these tools to simulate how humans would respond to questions or stimuli in different contexts (aher_using_2023; argyle_out_2023; arora2024express; binz2025foundation; dillion_can_2023; goli2024frontiers; hewitt2024predicting; horton_large_2023; park_generative_2023; qin2024aiturk; toubia2025twin).
 If these LLM-simulated experiments accurately approximate real human behavior, the implications for both academics and practitioners are substantial.
 Academics could use LLMs for pilot experiments to pinpoint stimuli with significant impact, thus improving the efficiency of theory development and experimental design.
-In fact, several review articles have already been written about this recent yet fast-growing literature (Bail,, 2024; Hutson and Mastin,, 2023; Sarstedt et al.,, 2024; Simmons and Hare,, 2023).
+In fact, several review articles have already been written about this recent yet fast-growing literature (bail2024can; hutson2023guinea; sarstedt2024using; simmons2023large).
 Firms could leverage these realistic simulations to explore different ideas and strategies, thereby improving customer insight and product development and optimizing marketing mix variables.
-Accordingly, in the recent past we have witnessed a large influx of firms offering services leveraging LLMs for customer insights (e.g., Synthetic Users, Outset AI, Nexxt, Voxpopme, Evidenza, Expected Parrot, Meaningful, xPolls, Ipsos, and CivicSync).
-Many practical applications of LLM simulations involve exploring counterfactual scenarios where one factor is altered while others remain constant, hence requiring causal estimates of treatment effects.
+Accordingly, in the recent past we have witnessed a large influx of firms offering services leveraging LLMs for customer insights (e.g., Synthetic Users, Outset AI, Nexxt, Voxpopme, Evidenza, Expected Parrot, Meaningful, xPolls, Ipsos, Pyxis and Generation Lab).
+In this paper we focus specifically on leveraging LLMs to simulate counterfactual human behavior.
+Indeed, many practical applications of LLM simulations involve exploring counterfactual scenarios where one factor is altered while others remain constant.
 For example, what would happen to sales if prices were increased, how would consumers respond if a message were framed differently, or how would market dynamics shift with the introduction of a new product?
-While the success of using LLMs to simulate counterfactual depends on the size and quality of their training or validation data that can correct for potential bias (Ludwig et al.,, 2024; Wang et al.,, 2024), the success also critically depends on the design of the simulations.
-Although extensive literature on experimental design with human subjects has established best practices and universal standards—such as random assignment and blinding—it remains unclear what constitutes best practices in designing LLM simulations and how these might differ from traditional experimental design principles.
+Early evidence on LLMs’ ability to simulate counterfactual human behavior is encouraging but mixed.
+For example, across 70 nationally-representative U.S.
+studies, hewitt2024predicting find that the correlation between predicted and observed treatment effect size is large ($r=0.85$), but the raw predictions derived from GPT-4 systematically overestimated actual effects, leading to a large (relative to human forecasters) RMSE between predicted and observed effects.
+Across 11 behavioral economics experiments, toubia2025twin find that LLM simulations achieve high predictive accuracy at the individual level, but replicated only about half of the treatment effects observed at the aggregate level.
+The ability of LLMs to simulate counterfactuals will likely improve as more powerful models are developed and the size and quality of the training and validation data continues to increase (ludwig2024large; wang2024large).
+In this paper, we focus on another key factor, which may be in fact much easier and cheaper to address and which is under the control of any user: the design of LLM experiments.
+We identify a fundamental issue that arises when simply applying best practices from human experiments to LLM experiments.
+This issue is not specific to a particular model or training dataset and, if left unaddressed, it is likely to cause some amount of residual bias in LLM simulations even as accuracy improves with better models and data.
+At the same time, we find this issue can be alleviated by carefully adjusting experimental best practices to the world of LLMs.
+In particular, many LLM simulations to date have relied on blinded experimental designs, where simulated participants are not aware of the experimental conditions or hypotheses being tested (argyle_out_2023; brand_using_2023; hewitt2024predicting; horton_large_2023; toubia2025twin).
+Blinding [^1] Also sometimes referred to as “masking” (american2019publication).
+has many benefits in human subject research: it reduces demand effects where participants modify their behavior based on what they believe researchers want to see, and it helps maintain ecological validity by presenting scenarios that more closely mirror real-world decision contexts.
+Following this convention, researchers typically construct LLM prompts that present a single scenario without revealing its experimental nature, similar to how human participants would experience the treatment condition in a between-subjects design.
+Although blinding is considered a best practice in human-subject experiments, we show that it can have unintended consequences in LLM simulations, which may no longer justify its benefits.
+Unlike human subjects and contexts that exist independently of the experiment, LLM-simulated subjects and contexts are ”created on the fly” based on prompts.
+When researchers make LLM-simulated “subjects” blind to experimental conditions, variations in the treatment can systematically affect unspecified variables that should remain constant, violating the unconfoundedness assumption typically required for valid causal inference.
+We use demand estimation as an example to empirically illustrate this challenge.
+We conduct experiments simulating consumer demand for 40 different products at different price points.
+We find that altering the product’s price not only affects the purchase outcomes but also systematically influences other variables that should not be influenced by the treatment—such as pre-treatment variables (e.g., demographic characteristics and customer purchase history) and contextual factors (e.g., price of competing products).
+This unintended variation in unspecified covariates violates the unconfoundedness assumptions, and in our case it also generates implausible demand curves when compared against an actual experiment with a representative sample of human participants.
+A tempting solution is specifying more variables in the prompts to address this confounding issue.
+However, we find this approach may sacrifice ecological validity because including more covariates can make the scenarios presented to the LLM less reflective of real-world decision-making contexts.
+For instance, when competing product information is explicitly mentioned in prompts, these variables become artificially salient in the LLM’s decision process, unlike typical consumers who may not always consider such information during routine purchases.
+This focalism issue creates a fundamental tension between controlling for confounding factors and maintaining ecological validity - a tradeoff that typically does not exist in human-subject experiments where covariates can be measured and controlled for during data analysis without making them artificially salient to participants during data generation.
+We present a theoretical framework illustrating that confoundedness in blind LLM simulations stems from prompts being ambiguous.
+To overcome this fundamental limitation and improve the simulation results, researchers should consider relaxing the blinding requirement and informing the LLM of the experimental design.
+This approach is not subject to the same confounding issues and holds promise for better generalizable performance as LLMs continue to advance.
+More broadly, we aim to contribute to the growing literature on LLM simulations by helping establish best practices in their design.
+Although extensive literature on experimental design with human subjects has established best practices and universal standards—such as random assignment and blinding—it remains unclear what constitutes best practices in designing LLM simulations and how these might differ from human-subject experimental design principles.
 Establishing LLM experimental design principles is crucial for enhancing the reliability and validity of LLM simulation results, thereby making them more valuable for both researchers and practitioners.
-Without clear guidelines on which experimental designs to adopt or avoid, the vast flexibility and sensitivity in LLM simulations can lead to widely varying outcomes (Carlson and Burbano,, 2024; Gao et al.,, 2024; Ludwig et al.,, 2024; Mohammadi,, 2024).
+Without clear guidelines on which experimental designs to adopt or avoid, the vast flexibility and sensitivity in LLM simulations can lead to widely varying outcomes (carlson2024use; gao2024take; ludwig2024large; mohammadi2024wait).
 When different simulation approaches yield conflicting results, researchers lack a basis for determining how much weight to place on each finding.
@@
 Developing such design principle is essential for advancing the field and ensuring that LLM simulations effectively contribute to our understanding of human behavior.
-Many LLM simulations to date have relied on blinded experimental designs, where simulated participants are not aware of the experimental conditions or hypotheses being tested (Argyle et al.,, 2023; Brand et al.,, 2023; Horton,, 2023).
-Blinding [^1] Also sometimes referred to as “masking” (Association et al.,, 2019).
-has many benefits in human subject research: it reduces demand effects where participants modify their behavior based on what they believe researchers want to see and often helps maintain ecological validity by presenting scenarios that more closely mirror real-world decision contexts.
-Following this convention, researchers typically construct LLM prompts that present a single scenario without revealing its experimental nature, similar to how human participants would experience the treatment condition in a traditional between-subjects design.
-Although blinding is considered a best practice in traditional experiments, we show that it can have unintended consequences in LLM simulations which may no longer justify its benefits.
-Unlike human subjects and contexts that exist independently of the experiment, LLM-simulated subjects and contexts are ”created on the fly” based on the prompt.
-When researchers make LLM-simulated “subjects” blind to experimental conditions, variations in the treatment can systematically affect unspecified variables that should remain constant, violating the unconfoundedness assumption typically required for valid causal inference.
-We use demand estimation as an example to empirically illustrate this challenge.
-We conduct experiments simulating consumer demand at different price points.
-We find that altering the product’s price not only affects the purchase outcomes but also systematically influences other variables that should not be influenced by the treatment—such as pre-treatment variables (e.g., demographic characteristics and customer purchase history) and contextual factors (e.g., price of competing products).
-This unintended variation in unspecified covariates violates the unconfoundedness assumptions, and we show that it also generates implausible demand curves when compared against an actual experiment with a representative sample of human participants.
-A tempting solution is specifying more variables in the prompts to address this confounding issue.
-However, we find this approach may sacrifice ecological validity because including more covariates can make the scenarios presented to the LLM less reflective of real-world decision-making contexts.
-For instance, when competing product information is explicitly mentioned in prompts, these variables become artificially salient in the LLM’s decision process, unlike typical consumers who may not always consider such information during routine purchases.
-This focalism issue creates a fundamental tension between controlling for confounding factors and maintaining ecological validity - a tradeoff that typically does not exist in traditional experiments where covariates can be measured and controlled for during data analysis without making them artificially salient to participants during data generation.
-We present a theoretical framework illustrating that confoundedness in blind LLM simulations stems from prompts being ambiguous.
-To overcome this fundamental limitation and improve the simulation results, researchers should consider relaxing the blinding requirement and informing the LLM of the experimental design.
-This approach is not subject to the same confounding issues and holds promise for better generalizable performance as LLMs continue to advance.
-To complement this approach, an opportunity for future research is to develop LLMs that can better interpret and understand experimental designs as well as natural language descriptions of counterfactual scenarios.
-More broadly, we aim to contribute to the growing literature on LLM simulations by providing a starting point for establishing best practices in their design.
-By highlighting the trade-offs between blinding, unconfoundedness, and ecological validity, we offer insights that can help researchers and practitioners design more reliable and valid LLM-based experiments.
-Our work underscores the importance of adjusting traditional experimental design principles to the unique characteristics of LLMs, ensuring that these powerful tools can be effectively harnessed to simulate and understand human behavior.
+In particular, by highlighting the trade-offs between blinding, unconfoundedness, and ecological validity, we offer insights that can help researchers and practitioners design more reliable and valid LLM-based experiments.
+In doing so, our work underscores the importance of adjusting human-subject experimental design principles to the unique characteristics of LLMs, ensuring that these powerful tools can be effectively harnessed to simulate and understand human behavior.
 The rest of the paper is organized as follows.
-Section 2 demonstrates the unintended consequences of blinding empirically using demand estimation as an example.
+Section 2 empirically demonstrates the unintended consequences of blinding using demand estimation as an example.
 Section 3 explores the standard solution of controlling for covariates and highlights its unique challenges in the context of LLM simulations.
-Section 4 proves an impossibility theorem demonstrating this fundamental issue will persist for future models as long as ambiguous prompting strategies are used, and examines its implications for techniques like fine-tuning and RAG.
+Section 4 argues theoretically that this fundamental issue will persist for future models as long as ambiguous prompting strategies are used.
 Section 5 discusses promising directions for making prompts unambiguous through unblinding, and Section 6 concludes.
@@
 This section uses GPT4, demand estimation and one common LLM simulation design to illustrate how blinding creates unique challenges in LLM simulations that are absent in human experiments, and Section 4 generalizes the findings to other versions of LLMs, other applications, and other simulation designs.
-One common way to implement blinding in LLM simulations is to conduct between-prompt experiments, where each prompt represents a single experimental condition without revealing the broader experimental context (Argyle et al.,, 2023; Brand et al.,, 2023; Horton,, 2023).
+One common way to implement blinding in LLM simulations is to conduct between-prompt experiments, where each prompt represents a single experimental condition without revealing the broader experimental context (argyle_out_2023; brand_using_2023; hewitt2024predicting; horton_large_2023; toubia2025twin).
 Blinded between-prompt LLM experiments have been used to mimic both between-subject and within-subject experiments with human participants.
@@
 To mimic a within-subject conjoint analysis study, researchers might present different product profiles or choice sets to the LLM in separate prompts, with each prompt describing only the current product profile or choice set.
-While random assignment in traditional human experiments ensures pre-treatment variables are balanced across conditions and contextual variables remain constant (even when unobserved), LLM-simulated participants and environments are generated dynamically.
-When researchers leave details unspecified in blind LLM prompts, the model must make (implicit or explicit) assumptions about these unspecified variables, which can systematically vary with the treatment and introduce confounding factors.
+While random assignment in human-subject experiments ensures pre-treatment variables are balanced across conditions and contextual variables remain constant (even when unobserved), LLM-simulated participants and environments are generated dynamically based on prompts.
+When researchers leave details unspecified in blind LLM prompts, the model may make (implicit or explicit) assumptions about these unspecified variables, which can systematically vary with the treatment and introduce confounding factors.
 We illustrate these issues using demand estimation, a classical problem in economics and marketing, where researchers are often interested in understanding how changes in a product’s price affect demand while holding everything else constant.
-We select the top 40 Consumer Packaged Goods (CPG) categories according to DellaVigna and Gentzkow, (2019), and for each category we pick one of the top selling products according to Walmart.com.
+We select the top 40 Consumer Packaged Goods (CPG) categories according to dellavigna_uniform_2019, and for each category we pick one of the top selling products according to Walmart.com.
 We record the regular price of each product on Walmart.com around April, 2024.
-Table A.1 in the Appendix documents this set of products.
+Table A.1 in the Web Appendix documents this set of 40 products.
 For each product, we vary the price from 0 to 200% of the regular price, in 20% increments.
-This gives us 11 price points to test for each product ($0\%\}$ , 20 % , ..100 % , ..200 % } of the regular price).
+This gives us 11 price points to test for each product ($\{0,20\%,..100\%,..200\%\}$ of the regular price).
 Throughout the paper, for each prompt template and for each {category,product} pair described in Table A.1, we vary the price systematically across these 11 price points.
-We refer to “relative price” as the price relative to the regular price ($\{-1,-0.8,...,0,...,0.8,1\}$ where 00 is the regular price).
+We refer to “relative price” as the price relative to the regular price ($\{-1,-0.8,...,0,...,0.8,1\}$ where 0 is the regular price).
 We use the Open AI API and we “pin” the model to gpt-4o-mini-2024-07-18 (the latest available at the time of running the experiment) to make sure that the exact same model is used for all our simulations.
-We adjust the “System” part of the prompt to increase the probability of receiving a usable answer from the model (the “System” portion of the prompt, which can be manipulated on the API and which is set to “You are a helpful assistant” on ChatGPT, provides meta-instructions to GPT).
+We adjust the “System” part of the prompt to increase the probability of receiving a usable answer from the model (the “System” portion of the prompt, which can be manipulated on the API, provides meta-instructions to GPT).
 Unless noted otherwise, for each prompt that corresponds to a {category, product, price} combination, we set the temperature to 1 and generate 50 draws to calculate the average response.
 First, we elicit unspecified variables from the LLM.
-For each {category, product, price} combination, we submit a prompt that asks GPT to fill in the blank contextual or pre-treatment variables given the randomized price (Kurita et al.,, 2019; Nozza et al.,, 2022; Zhang et al.,, 2020).
+For each {category, product, price} combination, we submit a prompt that asks GPT to fill in the blank contextual or pre-treatment variables given the randomized price (kurita2019measuring; nozza2022pipelines; zhang2020hurtful).
 Prompt 2 gives an example of such template that asks the past price given the current randomized price.
 Prompt 1: Unintended correlation between price and past price System: You, AI, are a customer.
-Your task is to fill in the blanks __.
+Your task is to fill in the blanks $_$.
 Return the completed information in comma-separated values, without any extra text.
@@
 The product is currently priced at: {price}.
-Return example 1: XX.XX
+Return example 1: $X$.XX
 Figure 1 demonstrates positive correlations between the price of the focal product and several variables, including the price that the customer paid last time, the price of the competing product, and the expiration days of the product.
-The x-axis is price relative to the baseline price, allowing us to aggregate across products.333 Prompts B and B in Appendix B document the prompts used to elicit the correlations between randomized price and competing price/expiration date.
-Appendix C presents additional analysis demonstrating the broader scope of these unintended correlations across multiple variables.
-While these correlations are plausible in observational data, they are undesirable in an experiment where researchers are interested in understanding what happens when estimating the treatment effect of changing the price of the focal product while holding everything else constant.
+The x-axis is price relative to the baseline price, allowing us to aggregate across products.
+[^3] Prompts B and B in Web Appendix B document the prompts used to elicit the correlations between randomized price and competing price/expiration date.
+Web Appendix C presents additional analysis demonstrating the broader scope of these unintended correlations across multiple variables.
+While these correlations are plausible in observational data, they are undesirable in an experiment where researchers are interested in understanding what happens when changing the price of the focal product while holding everything else constant.
 Figure 1: Unintended correlation between price and past price, competing price, and expiration days
-Next, we explore whether these unintended correlations are associated with confounding and biased estimates of relevant quantities when the goal is estimating causal treatment effects.
-To that end, we run an actual survey on human subjects, use a similar prompt in an LLM simulation, and compare the results.
+Next, we explore whether these unintended correlations are associated with biased estimates of relevant quantities when the goal is simulate counterfactual human behavior.
+We run an actual survey on human subjects, using a similar prompt in an LLM simulation, and compare the results.
 The survey with human participants (pre-registered, see https://aspredicted.org/pqys-st6w.pdf) was developed using Qualtrics.
-Each respondent answered one purchase intention question per product, with prices randomly drawn for each product and the order of products randomized across respondent.
+Each respondent answered one purchase intention question for each of the 40 products (see Appendix A for the complete list of categories, products, and prices), with prices randomly drawn for each product and the order of products randomized across respondent.
 [^4] We acknowledge that one difference between the study run with humans vs.
@@
 We then use Prompt 2 to simulate potential purchase decisions at different price points (where again {category,product} and {price} were varied according to our experimental design, mirroring the experiment with humans), and for each corresponding prompt, we used gpt-4o-mini-2024-07-18, with temperature set to 1, to generate 50 draws to calculate the average purchase probability at each {category,product,price} combination.
-Figure 2 shows the average demand curves obtained from GPT and from human participants.
-While the curves elicited from human participants follow a classic downward-sloping demand curve, the GPT-simulated curves are different from the curves elicited from human participants and follow an inverted-U shape.
-Although it is understandable that it is challenging for GPT to predict demand when price equals 0 as such scenario is probably not well represented in its training data, the curve remains inverted-U shape even if we remove the first point corresponding to a price of zero.
+Figure 2 shows the average demand curves (averaged over the 40 products) obtained from GPT and from human participants.
+While the curves elicited from human participants follow a classic downward-sloping demand pattern, the GPT-simulated curves follow an inverted-U shape.
+Although it is understandable that it is challenging for GPT to predict demand when price equals 0 as such scenario is probably not well represented in its training data, the curve remains inverted-U shaped even if we remove the first point corresponding to a price of zero.
 Prompt 2: Ask Purchase System: You, AI, are a customer.
@@
 [“purchase” or “not purchase”] Return example: purchase
-Figure 2: Demand curve elicited from humans vs. LLM
-These findings are not surprising in hindsight: in real observational data, the price of the focal product should be informative of the price of the competing products, especially if we do not know anything about the context, the type of store, and how the price for a product is set.
+Figure 2: Demand curve elicited from humans vs. LLM, averaged over 40 products
+These findings are not surprising in hindsight: in real observational data, the price of the focal product should be informative of other factors such as the price of the competing products, especially if we do not know anything about the context, the type of store, and how the price for a product is set.
 An LLM trained on such data would also naturally make such association.
-This association can be realistic and useful if researchers intend to use LLMs to simulate competing prices based on the focal product price.
+This association can be realistic and useful if researchers intend to use LLMs to elicit the correlation between these other factors and the focal product price.
 However, this same association is problematic when researchers intend to estimate the causal impact of changing the focal price, while holding everything else constant, because it leads to endogeneity.
@@
 If sufficient variation remains after controlling for covariates, then researchers can identify the causal effect of interest.
-In LLM-simulated experiments, a natural way to control for covariates is adding detailed information about individuals and their environment in the prompt.
+In LLM-simulated experiments, a natural starting point to control for covariates is adding detailed information about individuals and their environment in the prompt.
 These covariates can range from demographic information, location, preferences, budget, etc.
-We discuss other approaches including RAG and fine-tuning in the Section 4.3.
-Argyle et al., (2023) introduced an approach for controlling for such covariates.
+argyle_out_2023 introduced an approach for controlling for such covariates.
 They advocate for running studies on “silicon samples” of digital personas, which in their case were described by demographic variables and political beliefs that were drawn to align with nationally representative samples.
@@
 However, it does not fully resolve the issue.
-Appendix D describes the detailed prompts and covariates used, which involve the creating of 500 personas for each {product,category} pair.
+Web Appendix D describes the detailed prompts and covariates used, which involve creating of 500 personas for each {product,category} pair.
 Figure 3: Demand curves elicited by LLM after controlling for demographic variables
-An important consideration is the stage at which covariates are controlled for.
-When working with human data, empirical researchers often control for product characteristics and rich user history to address endogeneity during data analysis, after the data is already collected.
-Hence analyzing data does not change the original data generating process that already took place.
-In contrast, adding covariates in the prompt in LLM simulation occurs at the data generation stage.
+A natural question is what constitutes a principled approach to controlling for covariates in LLM simulations and what potential pitfalls exist.
+When working with human observational data, researchers commonly employ the “kitchen-sink” approach, controlling for many observed pre-treatment variables because a rich set of covariates makes the conditional-independence assumption more credible by capturing confounders.
+This benefit is strongest with many pre-treatment variables and a large dataset that yields better empirical overlap.
+While LLM simulations seem well suited here—one can specify many pre-treatment covariates and simulate large samples—an important consideration is the stage at which covariates are controlled for.
+When working with human data, empirical researchers often control for covariates such as product characteristics and rich user history to address endogeneity during data analysis, after the data have already been collected.
+Hence, analyzing the data does not change the original data-generating process that has already taken place.
+In contrast, adding covariates via the prompt in an LLM simulation occurs at the data-generation stage.
 Consequently, the additional covariates can change the factors considered by a simulated individual when making decisions.
@@
 However, while adding contextual variables such as competing prices in the LLM simulation may prevent the LLM from using the price treatment to infer these contextual variables, adding them may also lead the LLM-simulated customer to put a heavier emphasis on these factors when making decisions.
-Figure 4 shows that the demand curve (elicited by simulating the choices of 500 personas) is no longer smooth after fixing such contextual variables, [^5] We fix the competing price to be the same as the regular price of the focal product.
+Figure 4 shows (in the specific example of Coca-Cola) that the demand curve (elicited by simulating the choices of 500 personas) is no longer smooth after fixing such contextual variables, [^5] We fix the competing price to be the same as the regular price of the focal product.
 and the simulated customers’ purchase decisions follow a step function: customers will purchase if and only if the price is lower than or equal to the competing price.
-Appendix E describes the detailed prompts.
+Web Appendix D.1 describes the detailed prompts.
 [^6] One possible way to fix the non-smoothness of the demand curve is to simulate a distribution of contextual variables, each with difference past and competing prices.
@@
 In a real-world scenario, we should expect that even in the absence of heterogeneity in context (e.g., in a given store in a given day), there should still exists heterogeneity in consumers, such that the demand curve would be smoother.
-Figure 4: Purchase probability of Coca-Cola when controlling for demographic variables and competing price
-Such phenomenon in which decision makers place more emphasis on certain aspects of a choice set or scenario than they would in real-world decisions is known in the consumer behavior literature as focalism (Schkade and Kahneman,, 1998; Wilson et al.,, 2000), and constitutes a threat to ecological validity.
-For example, in the context of conjoint analysis, Bedi and Reibstein, (2021) argue that when minor attributes are presented together with major ones, they tend to receive undue attention from respondents (due to focalism), leading to upward biased estimates of the importance of these minor features.
-While it has been documented that LLM is prone to focalism (Cheng et al.,, 2023), it is important to emphasize that focalism is not an issue that is unique to LLM experiments.
-While traditional experiments with humans also run the risk of failing to simulate realistic scenarios, the nature of this risk differs.
-When running experiments with humans, using a blind design with realistic scenarios typically ensures ecological validity without introducing concerns of confoundedness.
-On the other hand, with LLM simulations, a blind design raises confoundedness concerns.
-Adressing confoundedness by adding covariates requires adding covariates in the data generation stage of the study, which raises yet another concern, focalism, which in turn reduces ecological validity.
-In other words, LLM simulations pose a unique dilemma to researchers.
-Improving ecological validity may require leaving unspecified details in the prompt, which runs the risk of introducing confoundedness.
-On the other hand, including extensive detail may reduce confounding, but at the expense of ecological validity due to focalism.
-4 Theoretical Framework
-So far we have demonstrated one empirical example with one specific version of an LLM.
-A natural question is whether the issue we have uncovered applies generally to all LLMs.
-This section presents a theoretical framework demonstrating that the challenges identified in this paper stems from ambiguous prompts rather than fundamental limitations of the models themselves.
-Therefore, the solution to this specific challenge lies not only in developing more advanced LLMs, but also in developing clearer prompting strategies.
-4.1 Blinding Generates Ambiguous Prompts
-To demonstrate how ambiguity arises in LLM prompts, we begin by examining a simple prompt $p$ under a blinded simulation design that illustrates the fundamental problem:
-Prompt 3: A simple ambiguous prompt Please consider the following product category: {category}.
+Such scenario is implausible in the real world because consumers are often inattentive and also care about other dimensions of the product characteristics.
+However, in LLM simulation, controlling for competing price to address endogeneity may unintentionally make the controlled covariates artificially salient to the simulated consumer and change the decision environment.
+Figure 4: Purchase probability, controlling for demographics and competing price: Coca-Cola example.
+Such phenomenon in which decision makers place more emphasis on certain aspects of a choice set or scenario than they would in real-world decisions is known in the consumer behavior literature as focalism (schkade1998does; wilson2000focalism), and constitutes a threat to ecological validity.
+For example, in the context of conjoint analysis, bedi2021damaged argue that when minor attributes are presented together with major ones, they tend to receive undue attention from respondents (due to focalism), leading to upward biased estimates of the importance of these minor features.
+While it has been documented that LLMs are prone to focalism (cheng2023compost) and focalism is not an issue that is unique to LLM experiments, the nature of the risk differs as it introduces a new set of trade-off: introducing additional covariates in the simulation may improve unconfoundedness but sacrifice ecological validity.
+Therefore, unlike standard experimental analysis where controlling covariates typically improve the performance, in LLM simulation, adding more covariates may degrade the performance, because it render the simulated human behavior less realistic.
+To illustrate this trade-off empirically, we leverage a publicly available dataset from toubia2025twin collected through multiple survey waves containing diverse measures of consumer cognitive and psychological traits alongside demographic information.
+Importantly, toubia2025twin also replicated our pricing study on their panel, allowing us to relate the results of a study with the exact same design as ours to other variables collected on the same sample of consumers.
+We implement a sequential covariate addition procedure across twelve stages, beginning with a baseline of 14 demographic variables (age, education, income, and others), then systematically adding behavioral economics and psychological measures one at a time.
+This granular approach introduces individual covariates sequentially: tightwad-spendthrift score (stage 2), discount rate and present bias (stage 3), risk and loss aversion (stages 4-5), financial literacy and numeracy (stages 6-7), mental accounting (stage 8), decision-making styles (stages 9-10), environmental attitudes (stage 11), and the Big Five personality traits (stage 12), expanding from 14 to 30 total control variables.
+Figure 5 tracks the mean absolute error (MAE) between actual consumer choices in toubia2025twin’s replication of our study and LLM predictions as we gradually add covariates.
+Appendix E provides detailed results showing the non-monotonic effects at each stage, Table E.1 reports the additional covariates added at each stage and Prompt E documents the prompt templates used.
+In standard experimental analyses with human data, controlling for additional covariates typically improves model performance, provided the sample size is sufficient.
+However, our results reveal a non-monotonic pattern: initial covariate additions improve accuracy, but beyond a threshold, further additions degrade performance.
+This deterioration may occurs because explicitly specifying covariates in the prompt induces focalism, in which the LLM assigns disproportionate weight to variables that consumers would typically ignore in naturalistic decision contexts, thereby compromising ecological validity despite reducing confounding.
+Figure 5: Mean absolute error progression as more covariates are controlled in the simulation. The MAE is calculated by comparing the purchase probability for each (product, price) combination.
+4 Core Issue: Ambiguous Prompting Strategy
+So far we have empirically explored one context (demand estimation) with 40 products and one specific version of an LLM.
+A natural question is whether the issue we have uncovered applies generally to all LLMs and other experiment contexts.
+This section formalizes the key issue and shows that one key step to harness the full potential of LLM in simulating human behavior lies in avoiding ambiguous prompting strategies.
+Formally,
+Definition 4.1 (Prompting Strategy).
+A prompting strategy is a function:
+$s:\mathcal{Q}\rightarrow\mathcal{P},$
+that maps each question $q\in\mathcal{Q}$ to a specific prompt $p\in\mathcal{P}$.
+Definition 4.2 (Ambiguous Prompting Strategy).
+An ambiguous prompting strategy occurs when there exists distinct questions $q_{1}$ and $q_{2}$ that have different correct answers, but are mapped to the same prompt $p$: $s(q_{1})=s(q_{2})=p$ for some prompt $p\in\mathcal{P}$.
+Such ambiguous prompting strategies should be avoided because - for any given LLM model, no matter how advanced they are, if researchers use an ambiguous prompting strategy, then if it happens to correctly answer one question correctly, $q_{1}$, it will incorrectly answer its ambiguous counterpart, $q_{2}$.
+Web Appendix F.1 provides a more detailed proof.
+An important step for practitioners to avoid ambiguous prompting strategy is to examine whether a specific prompt being used is open to multiple interpretations.
+For example, consider a simple prompt $p$ under a blinded simulation design:
+Prompt 3: A simple ambiguous prompt Consider the product category: {category}.
 Suppose you are in a grocery store, and you see the following product in that category: {product}.
 The product is priced at: {price}.
-Would you or would you not purchase the product?
-The design being blind, incomplete information is provided about the distribution from which the various variables in the prompt are drawn.
-As a result, this prompt is ambiguous.
-It allows for at least two interpretations:
+You decide to ___ [’purchase’ or ’not purchase’]
+This prompt is ambiguous because it allows for at least two interpretations:
   • $q_{1}$ (interventional): what is the customer’s purchase decision when the price is set to be $\{price\}$?
   • $q_{2}$ (observational): what is the customer’s purchase decision when the price happens to be $\{price\}$?
-The answers to these two questions are likely to be different.
-Using the do-operator (Pearl,, 2009) to distinguish mere observations and active intervention, and using $Y$ to represent $Purchase$ and $D$ to represent Price, the answer to $q_{1}$ can be represented as $a_{1}:P(Y|do(D))$ while the answer to $q_{2}$ can be represented as $a_{2}:P(Y|D)$.
-Consider a true data-generating process (DGP) described by a linear structural equation:
-$Y=\beta_{M}D+U$
-where $\beta_{M}$ represents the causal effect of $D$ on $Y$ through a set of realistic mechanism $M$, and $U$ represents unobserved factors.
-In a well-designed experiment, we have $Cov(U,D)=0$, making it possible to identify $\beta_{M}$ and estimate $P(Y|do(D))$ from data on $(Y,D)$ alone.
-But under a blinded design with an ambiguous prompt, the LLM does not know it should mimic an experimental environment.
-That is, the LLM does not have enough information to draw $U$ from the appropriate distribution (literally or figuratively - we do not attempt to model how the LLM makes inference but rather present a conceptual model).
-If the model is mostly trained on observational data, it is mostly likely going to mimic observational data patterns, where $Cov(U,D)\neq 0$.
-As a result, the LLM’s responses will reflect confounded relationships rather than the intended causal effects.
-Figure 5 summarizes two possible DGPs mimicked by the LLM under this simple ambiguous prompt.
-$D$$M$$Y$$U$(a) DGP for Experimental $D$$M$$Y$$U$(b) DGP for Observational Data
-Figure 5: Possible DGPs mimicked by LLM simulations under ambiguous simple prompts
-As discussed in Section 3, one might attempt to address this issue by adding more variables $X$ to the prompt.
-In principle, if $Cov(U,D|X)=0$, then controlling for $X$ can help restore identification.
-However, simply specifying covariates in the LLM prompt does more than condition on them statistically.
-Including $X$ may cause the simulation to alter the underlying decision-making mechanisms $M$ and shift toward a new set of mechanisms $M^{\prime}$ and produce an environment that differs substantially from the original scenario – the estimated parameter $\beta_{M^{\prime}}$ may have limited ecological validity, since it no longer reflects the realistic mechanisms $M$ that were originally of interest (as we illustrated in Figure 4).
-Figure 6 summarizes two possible DGPs mimicked by the LLM under this more complicated ambiguous prompt.
-[…]
-Figure 6: Potential DGPs mimicked by LLM simulations when controlling for additional variables
-These graphs reveal a deeper challenge: adding variables to the prompt can actually increase rather than decrease ambiguity from a causal perspective.
+These two questions yield different answers.
+Using the do-operator (Pearl2009) to distinguish active intervention from observation, and denoting $Y$ as Purchase and $D$ as Price, the answer to $q_{1}$ is $a_{1}:P(Y|do(D))$, where $do(D)$ is randomized and independent of unobserved factors $U$ ($D\perp U$).
+The answer to $q_{2}$ is $a_{2}:P(Y|D)$, where $D$ can be correlated with other confounders $U$ that affect $Y$ ($D\not\perp U$), as in observational data.
+Since LLMs are trained primarily on observational data, it is plausible that an ambiguous question will also be interpreted in a way that is similar to observational data, hence answering the wrong question.
+Appendix F.2 provides a DAG illustration of the different types of data generating process that an LLM may mimic when an ambiguous prompt is used.
+As discussed in Section 3, one might attempt to address this issue by adding more variables to the prompt.
+However, adding variables to the prompt can actually increase rather than decrease ambiguity from a causal perspective.
 To illustrate this multiplication of possible interpretations in practice, consider the following detailed example:
@@
 Would you or would you not purchase the product?
-This prompt, while detailed, is even more ambiguous. It permits more valid causal interpretations:
+This prompt, while detailed, may be even more ambiguous. It permits more valid causal interpretations:
   1. 1.
-[…]
+    Price as Treatment: $P(Y|do(\text{price}),\text{health\_info},\text{sugar\_content})$
     “What happens when we experimentally set price to \$4.99, holding fixed that customers are health-aware?”
   2. 2.
-[…]
+    Health Information as Treatment: $P(Y|\text{price},do(\text{health\_info}),\text{sugar\_content})$: “What happens when we experimentally expose customers to health articles, holding fixed price at \$4.99?”
   3. 3.
-[…]
-4.2 Impossibility Theorem for Ambiguous Prompting Strategies
-Having demonstrated how ambiguity can arise in practice, we now formalize these insights into a theoretical framework that explains why such ambiguity poses fundamental challenges for LLM evaluation and development.
-Definition 4.1 (Functional Representation of a Large Language Model).
-A large language model may be represented by:
-$f:\mathcal{P}\rightarrow\Delta(\mathcal{Y}),$
-where $\mathcal{P}$ is the set of all possible prompts, $\mathcal{Y}$ is the set of all possible outputs or responses, and $\Delta(\mathcal{Y})$ represents the set of all probability distributions over $\mathcal{Y}$.
-For any given prompt $p\in\mathcal{P}$, the model $f$ outputs a probability distribution $f(p)$ over potential responses $y\in\mathcal{Y}$.
-[…]
-Definition 4.2 (Prompting Strategy).
-A prompting strategy is a function:
-$s:\mathcal{Q}\rightarrow\mathcal{P},$
-that maps each question $q\in\mathcal{Q}$ to a specific prompt $p\in\mathcal{P}$.
-While prompting strategies can be refined empirically using validation datasets, developing a theoretical framework is essential for understanding when and why these strategies will generalize beyond the validation set.
-When ground truth data is available, researchers can experiment with different prompts and even fine-tune LLMs to improve performance on specific questions.
-However, without theoretical guidance, such empirical refinements may not yield strategies that reliably generalize to novel scenarios.
-This raises a fundamental question: what characteristics make a prompting strategy robust and reliable for out-of-distribution inference?
-We identify a particular class of prompting strategies to avoid:
-Definition 4.3 (Ambiguous Prompting Strategy).
-An ambiguous prompting strategy occurs when there exists distinct questions $q_{1}$ and $q_{2}$ that have different correct answers, but are mapped to the same prompt $p$: $s(q_{1})=s(q_{2})=p$ for some prompt $p\in\mathcal{P}$.
-Such ambiguous prompting strategies should be avoided because no LLM models, no matter how advanced they are, can answer all questions correctly given such prompting strategy:
-Theorem 4.1.
-[Impossibility Theorem for Ambiguous Prompting Strategies] For a given ambiguous prompting strategy, there exists no LLM $f$ that can correctly answer all questions correctly.
-Proof.
-Assume that $s(q_{1})=s(q_{2})=p$ for two distinct questions $q_{1}$ and $q_{2}$ that have different correct answers $a_{1}$ and $a_{2}$, with $a_{1}\neq a_{2}$.
-If the model can correctly answer $q_{1}$, such that $f(s(q_{1}))=f(p)=a_{1}$, then the model cannot answer $q_{2}$ because $f(s(q_{2}))=f(p)=a_{1}\neq a_{2}$.
-∎
-4.3 Implications for model development using fine-tuning or RAG
-Theorem 4.1 reveals how ambiguous prompting strategies create limitations when developing and deploying LLMs, even with techniques like fine-tuning or retrieval-augmented generation (RAG).
-Suppose developers leverage additional training data $(q_{1},a_{1})$ with a prompting strategy where $p=s(q_{1})$.
-If this prompt is ambiguous such that $p=s(q_{2})$ for some other question $q_{2}$ with correct answer $a_{2}\neq a_{1}$, then any fine-tuning that optimizes the model’s accuracy on $q_{1}$ necessarily sacrifices its accuracy on $q_{2}$.
-This trade-off is especially problematic given uncertainty about future applications - users may interpret the same prompts in different ways which developers cannot anticipate during model development.
-Consider a concrete example where developers have access to an experimental dataset that randomizes Coca-Cola’s price while keeping Pepsi’s price at business-as-usual levels.
-If developers fine-tune an LLM using a blinded design that simply describes purchase decisions at different Coca-Cola prices without revealing their experimental nature, they create an ambiguous prompt that permits both experimental and observational interpretations.
-While the fine-tuned model may excel at predicting purchases in scenarios matching the training data ($q_{1}$), this capability is less valuable since the experimental data already directly answers such questions.
-The more important question is whether the model can learn the causal relationship between price and purchases, and can generalize to answer novel questions ($q_{2}$) - like what happens when Pepsi’s price changes while Coca-Cola’s price remains at business-as-usual levels, or when both prices increase simultaneously.
-Such generalization proves challenging theoretically.
-First, under ambiguous prompts, fine-tuning based on experimental data provides no guarantee that the model will learn causal relationships.
-Suppose that the data generating process for the experimental dataset follows $Y=\beta_{M}D+U$, where $Y$ represents the purchase decision, $D$ represents the price of Coca-Cola, and $U$ represents everything else.
-The fine-tuned model may implicitly learn a regression coefficient $\widehat{b}=\frac{Cov(Y,D)}{Var(D)}=\beta_{M}+\frac{Cov(U,D)}{Var(D)}$.
-However, since the fine-tuned model builds on the base model, it can maintain a prior that $\widehat{\frac{Cov(U,D)}{Var(D)}}=\rho\neq 0$ and hence will not interpret the regression coefficient $\widehat{b}$ as causal.
-After accounting for this prior, the model may even incorrectly assume that the causal effect is $\widehat{\beta}_{M}=\widehat{b}-\rho$.
-Furthermore, even if the model somehow updates its assumptions correctly to handle one specific type of question ($q_{1}$), it may fail to generalize to other closely related but distinct questions ($q_{2}$) where different assumptions are needed.
-For example, suppose the model learns from experimental data where Coca-Cola’s price was randomized while Pepsi’s price remained business-as-usual.
-Through fine-tuning, the model might correctly learn to treat Coca-Cola’s price as independent of other unobserved market factors when analyzing that specific scenario ($q_{1}$).
-However, when asked to predict what happens when Pepsi’s price is set to another level while Coca-Cola’s price remain business-as-usual ($q_{2}$), the model might incorrectly maintain this assumption of independence, when in fact Coca-Cola’s business-as-usual price would likely be correlated with other market factors in this new scenario.
-This limitation is avoidable if the experimental nature of the data is explicitly communicated to the model during both training and evaluation through less ambiguous prompts.
-While the above example focuses on a case with one experimental dataset, the challenge persists even when using diverse experimental datasets for fine-tuning or retrieval-augmented generation (RAG).
-With RAG, this challenge manifests at both the development and deployment stages.
-During model development, the reference datasets should be annotated with metadata specifying which variables were randomized treatments versus correlated covariates.
-During retrieval, prompts should unambiguously specify the causal question to ensure the model draws from experiments with matching causal structures rather than just similar surface features.
-Without such explicit specification of the causal framework, neither fine-tuning nor RAG can reliably learn to distinguish between different types of variation in the data.
-To summarize, our theoretical framework proves that no general-purpose model can accurately simulate all counterfactual scenarios under an ambiguous prompting strategy.
-This statement should not be interpreted as suggesting that fine-tuning or RAG are not valuable tools - these techniques can substantially enhance model performance by incorporating high-quality training data and reducing various forms of bias.
-Rather, it demonstrates that while better training data and advanced techniques are important for improving simulation accuracy, they are complementary to, rather than substitutes for, clear and unambiguous prompt design.
-While obtaining better data for validation, fine-tuning, or bias-correction can be resource-intensive, implementing unambiguous simulation designs may be achieved at no additional cost.
-The optimal approach combines both strategies: leveraging advanced training techniques while ensuring an unambiguous prompting strategy.
-5 Unblinding as a Potential Solution
-Our theoretical framework points toward a promising direction: designing unambiguous prompts by explicitly communicating the experimental design in the prompt, which by definition entails unblinding.
-Here we discuss several possible simulation designs that leverage unblinding.
-One approach maintains the between-prompt structure where each prompt simulates one experimental condition, but explicitly reveals the full experimental design in each prompt.
-This allows generating multiple responses to form a distribution for each condition while maintaining the independence of observations across prompts.
-Another possibility is to use a within-prompt design where multiple experimental conditions are presented together in a single prompt with full information on the experimental design.
-This approach is more efficient computationally since it obtains responses for multiple conditions in one call to the model.
-However, showing multiple conditions together may induce artificial consistency in the responses.
-[^8] It is worth noting that within-prompt designs are not necessarily unblinded - a prompt could present multiple conditions (e.g., different prices) without explaining how those conditions were selected or distributed.
-Therefore, merely switching to within-prompt experiments is not sufficient to resolve prompt ambiguity.
-An even simpler approach is an unblind within-prompt aggregate design that directly asks the LLM to predict the aggregate results of the full experiment.
-This represents the greatest departure from traditional experimental protocols with human subjects.
-Comparing the relative performance of these different approaches is beyond the scope of this paper.
-Unlike our analysis of blinding - where we demonstrated fundamental limitations that will persist across model generations - the effectiveness of different unblinded designs may vary substantially across models and contexts.
-The relative performance one may find now may no longer hold in the future, especially if future models are trained or fine-tuned based on these prompts.
-Due to this reason, this section only focuses on demonstrating the feasibility of the unblinded design through the simplest approach of directly eliciting aggregate predictions of the results of the experiment.
-For this demonstration, we directly ask the LLM to predict aggregate experimental results across randomly assigned prices.
-Figure 7 shows that while this strategy’s predictions still deviate from actual consumer behavior, it does outperform the other potential solutions tested in this paper.
-Appendix F documents the detailed prompts.
-Figure 7: Unblinded design asking directly for aggregate predictions
-This approach is by no means perfect.
-In addition to the fact that the result still deviates from the ground truth, Figure F.1 shows that the results are also sensitive to the experimental range provided in the prompt: when the demand curve is built by combining two experiments that vary prices from [0%, 100%] and from [100%, 200%] of the regular price respectively, rather than a single experiment with prices varying from [0%, 200%], the estimates change.
-There are several possible reasons for this.
-One is that this LLM is unable to correctly process experimental information.
-Another possibility is that the existing prompt still has some remaining ambiguity.
-For example, the LLM may make assumptions about the setting, such as the type of stores and the customers, based on the set of prices included in the experiment.
-Appendix F conducts additional analysis which suggests that this sensitivity to the experimental range is partly driven by the current LLMs not being able to correctly process experimental design information, and shows that newer LLMs demonstrate improved (though still imperfect) consistency across different experimental ranges.
-Therefore, unlike the fundamental limitations of blinding that future LLMs may not easily address, sensitivity to experimental design information could potentially be resolved through model improvements, making this a promising area for future research.
-Figure 8: Unblinded design asking directly for aggregate predictions: price range comparison
+    Sugar Content as Treatment: $P(Y|\text{price},\text{health\_info},do(\text{sugar\_content}))$: “What happens when we experimentally highlight sugar content, holding fixed price and prior health awareness?”
+5 The value of unblinding
+Our theoretical framework points toward a promising direction: designing unambiguous prompts by explicitly communicating the experimental design, which by definition entails unblinding.
+The causal inference literature (AngristPischke2009; ImbensRubin2015; athey2017econometrics) has long emphasized the importance of clearly specifying estimands and sources of variation.
+Following this principle, we argue that unblinded prompts should specify both the treatment being randomized and the distribution from which the treatment is drawn.
+We test unblinding through two approaches: first by examining out-of-the-box LLMs to establish baseline differences, and then by generalizing to fine-tuning experiments.
+We begin by examining whether unblinding improves performance in standard, pre-trained language models, establishing a baseline for how prompt design affects predictions without task-specific training.
+To assess the robustness of the effects, we evaluate a series of GPT models spanning multiple generations, from GPT-3.5-Turbo to the latest models available at the time of writing (GPT-4.1).
+We operationalize the distinction between blinded and unblinded approaches through two different system prompts.
+Prompt 5: Blinded System Prompt System: You, AI, are a customer. Your task is to fill in the blank ___. Return the completed information without extra text.
+The unblinded condition alters the system prompt to include experimental context:
+Prompt 6: Unblinded System Prompt System: You, AI, are an expert in predicting customer behavior.
+The customer is given a survey on their purchase decision for the {product} in {category} where the price of the product is randomly and uniformly drawn from {min_price} to {max_price}.
+The customer is only presented with one price and is blind to this randomization design.
+The customer is given the following survey.
+Your task is to fill in the blank ___.
+Return the completed information without extra text.
+This design choice operationalizes our theoretical framework.
+The unblinded prompt explicitly communicates the data-generating process, including both the randomization mechanism and the support of the price distribution.
+Figure 6 reveals a consistent pattern across model generations.
+Unblinded prompts systematically generate more accurate predictions, with performance improvements ranging from 1% to over 60% relative to baseline.
+The mean absolute error (MAE) reduction persists across all models tested, including reasoning models, and this performance advantage remains substantial as models advance.
+Figure 6: Gain from unblinding for various models (ordered by release date and reasoning ability).
+The figure shows MAE performance improvement when moving from blinded to unblinded prompts, with GPT models ordered by release date.
+The MAE is calculated by comparing the purchase probability for each (product, price) combination, and then the improvement is normalized by the baseline MAE of the blinded prompts.
+Positive values indicate superior performance for unblinded prompts.
+While these out-of-the-box results demonstrate the value of unblinding, they raise a natural question: will this technique become obsolete as LLMs improve, particularly through fine-tuning?
+We argue that unblinding is not merely a strategy for evaluating and using pre-trained models but an important, complementary step in the model development and fine-tuning process itself.
+To illustrate, we consider fine-tuning a model on actual human data.
+When performing fine-tuning, researchers must decide what context to provide.
+For instance, should they reveal the experimental origins of the data (an unblinded approach, as in Prompt 5) or choose to not disclose such context (a blinded approach, as in Prompt 5)?
+This choice is consequential.
+Even if fine-tuning helps the model learn the correlation between price and purchase, without the context of randomization, it has an insufficient basis for interpreting this correlation as causal.
+Explicitly stating that the price was randomized provides context; hence, it may help the model interpret the learned relationship as causal and thereby improving its potential to generalize.
+Appendix F.3 provides a more detailed discussion.
+To empirically demonstrate this complementarity, we fine-tune models using our demand estimation survey data.
+We design an out-of-distribution test by performing a leave-one-category-out cross-validation.
+We divide the 40 products into four category groups: beverages (including juices, soft drinks, beer, and dairy drinks), refrigerated goods (including meats, dairy products, and frozen foods), snacks and bakery (including chips, cookies, coffee, candy, and ice cream), and household and pet supplies (including cleaning products, paper goods, pet food, and medications).
+In each fold, we fine-tune a model on three category groups and validate on the held-out fourth category group.
+This approach tests for true generalization, as the model must learn principles that transfer across fundamentally different product types rather than simply interpolating between similar items.
+This mirrors real-world applications where practitioners may be interested in forecasting scenarios that historical data has not observed before.
+For each of the four cross-validation folds, we fine-tune GPT-4o-mini models under a 2 (blinded vs.
+unblinded prompt in fine-tuning stage) × 2 (blinded vs.
+unblinded prompt in evaluation stage) factorial design, varying the prompt type (blinded Prompt 5 vs.
+unblinded Prompt 5) used during both the fine-tuning and evaluation stages.
+This design allows us to disentangle the effects of unblinding at each stage.
+To evaluate performance, we test each model on its held-out category, generating 50 responses per product-price pair (with temperature set to 1) and calculating the purchase probability.
+We then measure the MAE against the true probabilities from our human survey.
+Table 1 shows that fine-tuning consistently improves performance over the out-of-the-box model.
+However, the most accurate results are achieved when the model is both fine-tuned and evaluated with unblinded prompts.
+This combination reduces the MAE by an additional 15% compared to the best-performing blinded model, from 0.134 to 0.113.
+This finding highlights that unblinding offers a significant, nearly costless performance improvement, complementary to the more resource-intensive process of data acquisition and fine-tuning.
+Table 1: MAE of Fine-Tuned Models on Survey Data
+Model Evaluation Method
+Blinded Unblinded
+Out-of-box GPT-4o-mini 0.532 0.397
+Fine-tuned on Survey (Blinded) 0.134 0.128
+Fine-tuned on Survey (Unblinded) 0.130 0.113
+The value of unblinding becomes even more pronounced when fine-tuning models on complex or mixed-quality datasets.
+While the previous exercise demonstrated unblinding’s value with a single-source high-quality dataset, practical applications often involve greater uncertainty.
+Developers may know a model’s general purpose—such as forecasting demand under different pricing strategies—without knowing its exact future applications.
+Even in a simple setting where a firm only sells two products A and B, the application can include 1) forecasting demand when A and B are priced as usual 2) adjusting the price for product A when B is priced as usual 3) adjusting the price for product B when A is priced as usual, 4) adjusting price for both product A and B, 5) predicting the introduction of a new product C when product A and B are priced as usual.
+To address this range of potential queries, models may be fine-tuned on diverse training data comprising both experimental datasets, each with specific treatments and covariates, and observational datasets that capture business-as-usual patterns and correlations.
+While observational data can answer routine questions, it may prove problematic for certain counterfactual analyses.
+Without additional context during fine-tuning, experimental and observational data risk becoming indistinguishable to the model.
+Moreover, ambiguous prompts leave the model uncertain about which dataset to leverage for specific questions.
+Unblinding thus enhances model robustness against irrelevant datasets that may not directly address researchers’ intended questions.
+To illustrate the point, we replicate the aforementioned fine-tuning exercise, with the exception that we also added an irrelevant observational dataset that has documented how customers browse online on Amazon (berke2024open).
+The dataset is by construction observational and biased because it has only recorded purchases and no recorded non-purchases.
+Appendix G documents the exact prompt we used to incorporate the additional Amazon data in the fine-tuning process.
+Table 2 shows that when this irrelevant dataset is included, the performance of models trained with blinded prompts degrades significantly.
+In contrast, models trained with unblinded prompts prove far more robust.
+By providing clear context about the data’s origin, unblinding enables the model to correctly weight the experimental data and discount the noisy observational data, preserving its predictive accuracy.
+Table 2: MAE of Fine-Tuned Models on Mixed (Survey + Amazon) Data
+Model Evaluation Method
+Blinded Unblinded
+Out-of-box GPT-4o-mini 0.532 0.397
+Fine-tuned on Survey+Amazon (Blinded) 0.233 0.126
+Fine-tuned on Survey+Amazon (Unblinded) 0.145 0.120
 6 Conclusion
@@
 This confounding makes it difficult to accurately simulate counterfactual human behavior.
-More generally and fundamentally, making the LLM blind to the experimental design leads to ambiguous prompts.
-Given LLMs are trained primarily on observational data, ambiguous prompts tend to favor correlational rather than causal observations, making variables unspecified in the prompt susceptible to the intended intervention.
-Because these challenges stem from ambiguous prompting strategies rather than model limitations, our theoretical framework suggests that it is not possible to fully resolve it through improved training data.
+Our exploration of potential solutions to this challenge reveals important insights for improving LLM simulations.
+While adding detailed context to control for confounders can help address the confounding issue, our analysis demonstrates that this approach can introduce focalism—causing simulated individuals to overweight explicitly mentioned variables relative to realistic human behavior.
+This trade-off between reducing confounding and maintaining ecological validity represents a fundamental consideration when designing LLM simulations with detailed prompts.
+These empirical observations point to a deeper theoretical issue.
+Making the LLM blind to the experimental design leads to ambiguous prompts.
+Because these challenges stem from ambiguous prompting strategies rather than model limitations, our theoretical framework suggests that developing unambiguous prompting strategies constitutes a fundamental principle that complements other model improvement techniques.
+Building on this theoretical understanding, our empirical findings demonstrate that unblinded experimental designs offer a particularly promising path forward.
+By explicitly communicating the randomization mechanism and experimental context, unblinded prompts reduce ambiguity and enable more accurate simulations.
+This approach shows consistent improvements across model generations and proves especially valuable when combined with fine-tuning, where it enhances the model’s ability to distinguish causal from correlational patterns in training data.
+The complementary nature of unblinding and fine-tuning suggests that developing LLMs capable of accurately simulating counterfactual human behavior when provided with clear, unambiguous prompts represents both a crucial and achievable direction for advancing the field.
 Our findings have important practical implications for different stakeholders in the LLM ecosystem.
@@
 When prompts permit multiple interpretations, improving performance on one interpretation may degrade performance on others - an unnecessary trade-off that could be avoided through clearer prompting strategies, especially given uncertainty about future applications.
-We explored several potential solutions to this challenge.
-Adding detailed context to control for confounders showed promise but tends to introduce focalism which could invalidate results.
-A more promising approach consists in unblinded experimental designs that explicitly communicate randomization.
-This approach showed potential for reducing ambiguity, though its effectiveness may vary across models and contexts.
-Therefore, developing LLMs that can better follow the instruction to simulate counterfactual human behavior given unambiguous prompts can be a crucial direction for future research.
-Our paper highlights a broader insight of using LLMs: rather than simply applying human subject protocols to LLMs, researchers need to develop new experimental practices tailored to LLMs’ unique properties.
-The field of LLM simulations is still emerging, and best practices remain to be established.
-Our work provides a theoretical foundation for understanding one of the key challenges and evaluating potential solutions.
-We hope this framework will guide future research in developing more reliable methods for leveraging LLMs in experimental research, while remaining mindful of their fundamental differences from human subjects.
+Our paper highlights a broader insight regarding the use of LLMs: rather than simply applying human subject protocols to LLMs, researchers must develop new experimental practices tailored to LLMs’ unique properties.
+The field of LLM simulations remains nascent, and best practices have yet to be established.
+Our work highlights the need for a new set of principles for designing LLM simulations and evaluation frameworks.
+We hope this will guide future research in developing more reliable methods for leveraging LLMs in experimental research while remaining mindful of their fundamental differences from human subjects.
+Funding and Competing Interests Declarations
+All authors certify that they have no affiliations with or involvement in any organization or entity with any financial interest or non-financial interest in the subject matter or materials discussed in this manuscript.
+The authors have no funding to report.
 Appendix A Category, Product and Price Information
@@
 Cold Remedies -Adult Equate Value Size Honey Lemon Cough Drops with Menthol, 160 Count 4.68
-Note: The 40 categories were taken from DellaVigna and Gentzkow, (2019) and the top products and their regular prices from Walmart.com as of April, 2024.
+Note: The 40 categories were taken from dellavigna_uniform_2019 and the top products and their regular prices from Walmart.com as of April, 2024.
 Due to the dynamic nature of retail pricing, actual prices may vary.
@@
 Prompts B and B document the prompts used to elicit the correlations between randomized price and competing price/expiration days shown in Figure 1.
-Prompt 5: Correlation between price and competing price System: You, AI, are a customer.
-Your task is to fill in the blanks __.
+Prompt 7: Correlation between price and competing price System: You, AI, are a customer.
+Your task is to fill in the blanks $_$.
 Return the completed information in comma-separated values, without any extra text.
@@
 The price of a similar competing product from a different brand is      [a number with up to 2 decimal points].
-Return example 1: XX.XX
-Prompt 6: Correlation between price and expiration days System: You, AI, are a customer.
-Your task is to fill in the blanks __.
+Return example 1: $X$.XX
+Prompt 8: Correlation between price and expiration days System: You, AI, are a customer.
+Your task is to fill in the blanks $_$.
 Return the completed information in comma-separated values, without any extra text.
@@
 Appendix C Covariance of unspecified variables
-This section demonstrates how prevalent variables covary with each other. We consider a more detailed prompt C.1 that elicits demographic, contextual, and pre-treatment variables.
-Prompt 7: Correlation between multiple variables System: You, AI, are a customer.
+This section demonstrates how prevalent variables covary with each other, for which we use a more detailed prompt C.2 that elicits demographic, contextual, and pre-treatment variables.
+Figure C.1 summarizes the covariance matrix between these variables.
+On one hand, this covariance matrix demonstrates that the LLM is capable of simulating many realistic correlations between covariates.
+For example, household income is positively correlated with budget.
+However, several covariates are correlated with the price change variable.
+Specifically, the focal price is correlated with product characteristics such as expiration days of the product, contextual variables such as competing prices, and pre-treatment variables such as purchase history.
+In an experiment with humans, these variables would either be held constant or be randomly distributed across treatment conditions in a way that is orthogonal to the treatment.
+Figure C.1: Covariance matrix of demographic, contextual, and pre-treatment variables
+Prompt 9: Correlation between multiple variables System: You, AI, are a customer.
 Your task is to fill in the blanks  .
@@
 Figure C.2: Eliciting correlation between relevant variables
-Figure C.2 summarizes the covariance matrix between these variables.
-On one hand, this covariance matrix demonstrates that the LLM is capable of simulating many realistic correlations between covariates.
-For example, household income is positively correlated with budget.
-However, several covariates are correlated with the price change variable.
-Specifically, the focal price is correlated with product characteristics such as expiration days of the product, contextual variables such as competing prices, and pre-treatment variables such as purchase history.
-In an experiment with humans, these variables would either be held constant or be randomly distributed across treatment conditions in a way that is orthogonal to the treatment.
-Figure C.2: Covariance matrix of demographic, contextual, and pre-treatment variables
 Appendix D Simulation that controls for demographic variables
-We use the following prompt template to first generate the distribution of demographic variables and then control for demographic variables.
+We use Prompt D to first generate the distribution of demographic variables and then control for demographic variables.
 First, for each {category,product}, we leave the demographic variables as blank, and use the pinned version of gpt-4o-mini-2024-07-18 to generate 500 replications (personas) with a temperature of 1, giving us the joint distribution of demographic variables for that product.
@@
 Then we aggregate across all products at the same relative price to construct the aggregate demand curve.
-Prompt 8: Control for demographic variables System: You, AI, are a customer.
+Prompt 10: Control for demographic variables System: You, AI, are a customer.
 Your task is to fill in the blanks  .
@@
 [“purchase” or “not purchase”] Return example: 35, female, bachelor’s degree, 50000, software engineer, Asian,married, 3, 1, CA, own, 3.99, purchase
-Appendix E Simulation that controls for demographic variables + competing price
-To provide a concrete example, we consider Coca-cola and asks what happens to the purchase probability when the price of a competing product is set at a certain level.
-We follow the exact same two-step procedure as Appendix D, except that we have fixed the price of a competing product to be the same as the regular price of the focal product when eliciting purchase decisions from the 500 personas.
-Prompt 9: Control for demographic variables+ competing price System: You, AI, are a customer.
+D.1 Simulation that also controls for competing price
+To illustrate the unintended effect of adding covariates in simulation, we focus on a product, Coca-cola, and asks what happens to the purchase probability when the price of a competing product is set at a certain level.
+We follow the exact same two-step procedure as Web Appendix D, except that we have fixed the price of a competing product to be the same as the regular price of the focal product when eliciting purchase decisions from the 500 personas.
+Prompt 11: Control for demographic variables+ competing price System: You, AI, are a customer.
 Your task is to fill in the blanks  .
@@
 Different customers, each with their own characteristics, will come to the store and make their purchase decision.
-While it is plausible that some customers may notice the price of the competing product and use it as a reference point to make purchase decisions, not all customers are attentive and may be aware of the price of all other competing products.
-Therefore, it is implausible to assume that the purchase will be $100\%$ ($0\%$) when the price of the focal product is lower than or equal to (greater than) the price of competing products.
-Appendix F Direct simulation of the experimental results using unblinded design
-Prompt F is a prompt template we use to predict the experimental results using an unblinded design, where information about the experiment is directly provided.
-For each {category,product}, we replace {RegularPrice} with the corresponding price listed in Table A.1.
-For Figure 7, we replace the {Prices} with the set of prices that we experiment with, from 0% to 200% of regular price.
-For example, when the regular price is \$4.68.
-{Prices} are ”\$0.00, \$0.94, \$1.87, \$2.81, \$3.74, \$4.68, \$5.62, \$6.55, \$7.49, \$8.42, \$9.36”.
-In this basic version, we leave Additional Instruction 1 and 2 as empty strings “”.
-We set the temperature to 0 and use ’gpt-4o-mini-2024-07-18’.
-For Figure 7, in addition to considering the {Prices} to be from 0% to 200%, we also try a version from 0% to 100%, labeled as ”Lower_Half_Price_Range” and a version from 100% to 200%, labeled as ”Upper_Half_Price_Range”.
-It shows that the simulation result is sensitive to the range of prices provided in the experiment.
-To assess whether this sensitivity is driven by another form of prompt ambiguity, where LLMs may make unintended assumptions on the context given the set of experimental prices, in Figure F.1 we provide more explicit instructions by replacing {Additional Instruction 1} with ”Note that some elements of the experiments are chosen arbitrarily by the researchers and you should not rely on them to make your prediction.” and replacing {Additional Instruction 2} with ”Note that the Set of Prices to be experimented is chosen arbitrarily and you should not make your prediction based on this Set of Prices.” However, we find that the results are still sensitive, suggesting that the issue are driven by the current generation of LLMs that we are testing being unable to closely follow such instruction, rather than prompt ambiguity.
-To assess the potential for future models to make consistent prediction, we further test ’o1-mini-2024-09-12’ and ’o1-preview-2024-09-12’.
-Because they do not allow for ’System’ and only accept the ’User’ part of the prompt, we concatenate the ’System’ and ’User’ together as our input prompt.
-Figure F.1 shows that when using ’gpt-4o-mini-2024-07-18’, the inconsistency still remains substantial despite explicit instruction not to rely too much on the experimental range to make the prediction.
-However, ’o1-mini-2024-09-12’ and ’o1-preview-2024-09-12’ demonstrates increasingly decreased inconsistencies.
-[…]
-Figure F.1: Unblinded design with explicit instruction to disregard experimental range: Multiple consecutive model versions.
+While it is plausible that some customers may notice the price of the competing product and use it as a reference point to make purchase decisions, our step-wise non-smooth demand curve is implausible in real-world scenarios where not all customers are attentive or are aware of the price of all other competing products.
+Appendix E Non-monotonic impact of controlling for covariates
+Table E.1 gives an overview of the variables used and the corresponding results documented in Section 3.
+Table E.1: Sequential Covariate Addition in Sensitivity Analysis
+Chunks Total New New Covariates MAE
+Covariates Covariates (Names)
+1 14 14 Age, US Citizenship, Education Level, Employmen… 0.2250
+2 15 1 Tightwad-Spendthrift Score 0.0971
+3 17 2 Discount Rate, Present Bias 0.0989
+4 18 1 Risk Aversion 0.1413
+5 19 1 Loss Aversion 0.1478
+6 20 1 Financial Literacy 0.1524
+7 21 1 Numeracy 0.1714
+8 22 1 score mentalaccounting 0.1646
+9 23 1 score maximization 0.1700
+10 24 1 score minimalism 0.1842
+11 25 1 score GREEN 0.2025
+12 30 5 score agreeableness, score extraversion, score … 0.1929
+  • Note: This table presents the sequential addition of covariates in the sensitivity analysis.
+Stage 1 includes sociodemographic characteristics (age, citizen status, education, employment status, household size, marriage, political affiliation, political views, race, region, religion, religious attendance, sex, and total family income).
+Stages 2-11 add individual behavioral economics and psychological measures.
+Stage 12 adds Big Five personality traits (agreeableness, extraversion, neuroticism, openness, conscientiousness).
+Mean Absolute Error (MAE) represents prediction error at each stage.
+Prompt E shows the final template that includes all the covariates we tested at Stage 12.
+At stage 1, only the demographic variables (the first 14 items) were included.
+Each subsequent stage systematically added behavioral and psychological measures: tightwad-spendthrift (stage 2), temporal discounting (stage 3), risk aversion (stage 4), loss aversion (stage 5), financial literacy (stage 6), numeracy (stage 7), mental accounting (stage 8), maximization (stage 9), minimalism (stage 10), environmental attitudes (stage 11), and finally the Big Five personality traits (stage 12).
+Prompt 12: Prompt Templates Based on Detailed Covariates: Part I System: You, AI, are a customer.
+Your task is to fill in the blanks  .
+Return the completed information in comma-separated values, without any extra text.
+User: You are a consumer with the following characteristics: # Demographics:
+- Geographic region: {region}
+- Gender: {sex}
+- Age: {age}
+- Education level: {education}
+- Race: {race}
+- Citizen of the US: {citizen_status}
+- Marital status: {marriage}
+- Religion: {religion}
+- Religious attendance: {religious_attendance}
+- Political affiliation: {political_affiliation}
+- Income: {total_family_income}
+- Political views: {political_views}
+- Household size: {household_size}
+- Employment status: {employment_status}
+# Tightwad-Spendthrift: {score_ST-TW} ({pct_spendthrift} percentile)
+<<note: The score ranges from 4 to 26. Lower scores (4-11) indicate difficulty spending money, while higher scores (19-26) indicate difficulty controlling spending.>>
+# Discount, Present Bias:
+<<note: These are implied rates computed from your time-value of money preferences.
+Higher values of the discount rate imply greater impatience.
+Higher values of present bias imply greater departure from normative economic behavior.>>
+- Discount: {score_discount} ({pct_discount} percentile)
+- Present Bias: {score_presentbias} ({pct_presentbias} percentile)
+# Risk aversion: {score_riskaversion} ({pct_riskaversion} percentile)
+<<note: Higher scores indicate a greater tendency for risk aversion in a choice between a sure-amount and lottery payout.>>
+Prompt 12: Prompt Templates Based on Detailed Covariates: Part II # Loss aversion: {score_lossaversion} ({pct_lossaversion} percentile)
+<<note: Higher scores indicate a greater tendency for loss aversion in a choice between a sure-amount and a lottery payout.>>
+# Financial Literacy: {score_finliteracy} ({pct_finliteracy} percentile)
+<<note: The score ranges from 0 to 8, and a higher score indicates you correctly answered more questions related to general financial literacy.>>
+# Numeracy: {score_numeracy} ({pct_numeracy} percentile)
+<<note: The score ranges from 0 to 8, and a higher score indicates you correctly answered more questions related to numeracy.>>
+# Mental Accounting: {score_mentalaccounting} ({pct_mentalaccounting} percentile)
+<<note: The score ranges from 0 to 100 percent, and higher scores indicate a greater adherence to the principles of mental accounting proposed by Thaler: segregate gains, integrate losses, segregate a small gain from a large loss, and integrate a small loss with a large gain.>>
+# Maximization: {score_maximization} ({pct_maximization} percentile)
+<<note: The score ranges from 1 to 5, and higher scores indicate a tendency to optimize rather than satisfice when making decisions.>>
+# Minimalism: {score_minimalism} ({pct_minimalism} percentile)
+<<note: The score ranges from 1 to 5, and a higher score indicates a higher preference for minimalism.>>
+# GREEN: {score_GREEN} ({pct_green} percentile)
+<<note: The score ranges from 1 to 5, and higher scores indicate a higher affinity for environmentalism.>>
+# Big 5 Personality:
+<<note: Openness reflects curiosity and receptiveness to new experiences, Conscientiousness indicates self-discipline and goal-directed behavior, Extraversion measures sociability and assertiveness, Agreeableness reflects compassion and cooperativeness, and Neuroticism captures emotional instability and susceptibility to negative emotions.
+Each score ranges from 1 to 5, and a higher score indicates a greater display of the associated traits.>>
+- Extraversion: {score_extraversion} ({pct_extraversion} percentile)
+- Agreeableness: {score_agreeableness} ({pct_agreeableness} percentile)
+- Conscientiousness: {wave1_score_conscientiousness} ({pct_conscientiousness} percentile)
+- Openness: {score_openness} ({pct_openness} percentile)
+- Neuroticism: {score_neuroticism} ({pct_neuroticism} percentile)
+Appendix F Theoretical Framework
+F.1 Impossibility of Ambiguous Prompting
+Theorem F.1.
+[Impossibility Theorem for Ambiguous Prompting Strategies] For a given ambiguous prompting strategy, there exists no LLM $f$ that can correctly answer all questions correctly.
+Proof.
+Assume that $s(q_{1})=s(q_{2})=p$ for two distinct questions $q_{1}$ and $q_{2}$ that have different correct answers $a_{1}$ and $a_{2}$, with $a_{1}\neq a_{2}$.
+If the model can correctly answer $q_{1}$, such that $f(s(q_{1}))=f(p)=a_{1}$, then the model cannot answer $q_{2}$ because $f(s(q_{2}))=f(p)=a_{1}\neq a_{2}$.
+∎
+F.2 Illustration of the underlying DGP that LLM simulation mimics if prompts are interpreted differently
+Consider a true data-generating process (DGP) described by a linear structural equation:
+$Y=\beta_{M}D+U$
+where $\beta_{M}$ represents the causal effect of $D$ on $Y$ through a set of realistic mechanism $M$, and $U$ represents unobserved factors.
+In a well-designed experiment, we have $\text{Cov}(U,D)=0$, making it possible to identify $\beta_{M}$ and estimate $P(Y|do(D))$ from data on $(Y,D)$ alone.
+But under a blinded design with an ambiguous prompt, the LLM does not know it should mimic an experimental environment.
+That is, the LLM does not have enough information to draw $U$ from the appropriate distribution (literally or figuratively - we do not attempt to model how the LLM makes inference but rather present a conceptual model).
+If the model is mostly trained on observational data, it is mostly likely going to mimic observational data patterns, where $\text{Cov}(U,D)\neq 0$.
+As a result, the LLM’s responses will reflect confounded relationships rather than the intended causal effects.
+Figure F.1 summarizes two possible DGPs mimicked by the LLM under this simple ambiguous prompt.
+DDMMYYUU(a) DGP for Experimental DataDDMMYYUU(b) DGP for Observational Data
+Figure F.1: Possible DGPs mimicked by LLM simulations under ambiguous simple prompts
+As discussed in Section 3, one might attempt to address this issue by adding more variables $X$ to the prompt.
+In principle, if $\text{Cov}(U,D|X)=0$, then controlling for $X$ can help restore identification.
+However, simply specifying covariates in the LLM prompt does more than condition on them statistically.
+Including $X$ may cause the simulation to alter the underlying decision-making mechanisms $M$ and shift toward a new set of mechanisms $M^{\prime}$ and produce an environment that differs substantially from the original scenario – the estimated parameter $\beta_{M^{\prime}}$ may have limited ecological validity, since it no longer reflects the realistic mechanisms $M$ that were originally of interest (as we illustrated in Figure 4).
+Figure F.2 summarizes two possible DGPs mimicked by the LLM under this more complicated ambiguous prompt.
+DDMMYYUUXX(a) Adding $X$ may address confoundingDDMMYYUUXX(b) Adding $X$ may also sacrifice ecological validity
+Figure F.2: Potential DGPs mimicked by LLM simulations when controlling for additional variables
+F.3 Implications for Model Development
+The impossibility theorem extends to fine-tuning and RAG approaches.
+Consider experimental data where Coca-Cola prices were randomized while Pepsi prices remained fixed.
+Fine-tuning on this data with ambiguous prompts cannot guarantee the model learns the causal relationship $\beta_{M}$.
+The model may learn the regression coefficient $\hat{b}=\beta_{M}+\text{Cov}(U,D)/\text{Var}(D)$ while maintaining a prior that $\text{Cov}(U,D)\neq 0$, leading to biased causal inference.
+Moreover, even if the model correctly learns independence assumptions for the training scenario ($q_{1}$), it may fail to generalize appropriately to new scenarios ($q_{2}$).
+When asked about cases where Pepsi prices vary and Coca-Cola prices are fixed, the model might incorrectly maintain the independence assumption from training, when correlation with unobserved factors would actually exist.
+Therefore, while fine-tuning and RAG remain valuable for incorporating domain knowledge, they are complements to rather than substitutes for unambiguous prompt designs.
+Appendix G Fine-tuning with additional irrelevant data
+To test the robustness of different fine-tuning strategy when additional dataset is included, we sample $10,000$ observations from the Amazon dataset berke2024open and incorporate them into fine tuning using Prompt G).
+Prompt 13: Prompt for Amazon data fine-tuning System: You, AI, are a customer.
+Your task is to fill in the blank ___.
+Return the completed information without extra text.
+User: Consider the product category: {category}.
+Suppose you are shopping online, and you see the product {product}.
+The product is priced at \${price}.
+You decide to ___ [“purchase” or “not purchase”].
+Return example 1: purchase Return example 2: not purchase